### v1

# Conceptualization of this study, Methodology, Software

Abstract
Continual learning (CL) refers to the ability of an algorithm to continuously and incrementally acquire new knowledge from its environment while retaining previously learned information. A model trained on one data modality often fails when tested with a different modality. A straightforward approach might be to fuse the two modalities by concatenating their features and training the model on the fused data. However, this requires retraining the model from scratch each time it encounters a new domain. In this paper, we introduce a continual learning algorithm capable of incrementally learning different data modalities by leveraging both class-incremental and domain-incremental learning scenarios in an artificial environment where labeled data is scarce, yet non-iid (independent and identical distribution) unlabeled data from the environment is plentiful. The proposed algorithm is efficient and only requires storing prototypes for each class. We evaluate the algorithm’s effectiveness on a challenging custom multimodal dataset comprising of tactile data from a soft pneumatic gripper, and visual data from non-stationary images of objects extracted from video sequences. Additionally, we conduct an ablation study on the custom dataset and the Core50 dataset to highlight the contributions of different components of the algorithm. To further demonstrate the robustness of the algorithm, we perform a real-time experiment for object classification using the soft gripper and an external independent camera setup, all synchronized with the Robot Operating System (ROS) framework.
keywords:
[…]
1 Introduction
In recent years, machine learning models have matched or even exceeded human-level performance on tasks such as image classification Ahmad et al., object recognition Zou et al., natural language processing (NLP) Otter et al., playing games in simulation, etc. While these models excel in static tasks where the data distribution remains constant, they encounter difficulties in adapting to changing environments and need to restart the training process whenever new data is introduced. On the other hand, humans are adept at learning from dynamic environments, continuously acquiring, updating, and applying knowledge over time. We expect AI systems to adapt in a similar fashion.
Continual learning (CL) Wang et al., Lesort et al., also known as lifelong learning or incremental learning, is a machine learning subfield that focuses on progressively training models on a continuous stream of data to accumulate and retain knowledge over time. However, a significant obstacle for these artificial models is their tendency to forget previously acquired skills when exposed to new tasks or data distributions. This issue, known as catastrophic forgetting, often leads to a substantial drop in performance as the new information overwrites the old knowledge or alters the data distribution. A naive solution to this problem is to save all data, shuffle it, and retrain the model offline from scratch. While this method effectively addresses catastrophic forgetting, it is time-inefficient, suboptimal, and unsustainable for modern large-scale models that require significant computational power. Continual learning aims to develop methods that balance performance and efficiency more effectively. Robots operating in unstructured real world environments will encounter new tasks and challenges over time, necessitating capabilities that cannot be fully anticipated from the outset. These robots need to learn continually, acquiring new skills without forgetting previously learned ones.
A model trained in one domain will be ineffective when deployed in a completely different domain or used with a different data modality than it was originally trained on. One straightforward approach might be to use separate models for each data modality or to jointly train a single model on all modalities offline. However, this is not a feasible solution for agents with limited memory and computational resources that learns new domains through active exploration and interaction with its environment. Another challenge is the limited availability of labeled data for supervised incremental training. A large research effort is dedicated towards learning from unlabeled data, which is often more readily available than labeled data in practical applications. Within this effort, the field of semi-supervised learning (SSL) Van Engelen and Hoos has recently demonstrated the most promising results. In this setting, an agent can utilize a significant amount of unlabeled data from its environment. It can employ the knowledge gained from labeled data to learn new objects of the same class or to enrich its understanding of previously learned classes using the unlabeled data.
In this paper, we consider a realistic continual learning problem where a single continual learning model learns different modalities of data incrementally by leveraging both class-incremental and domain-incremental learning scenarios in an artificial environment where labeled data is scarce yet non-iid (independent and identical distribution) unlabeled data from the environment is plentiful. We mainly experiment with two different SSL conditions: Radom images, where for each object of the same parent class we randomly sample some images from the dataset that are not included in the training data; Unique objects, where we sample the images of new unseen objects that belong to the same parent class. More details about the SSL conditions is provided in section 3.3.
To perform a multimodal CL experiment, we leverage two different modalities of data mainly the tactile data coming from sensors integrated with a soft gripper and the visual data from an independent camera setup. The CL model is designed to learn the classes in both the domains incrementally. To address this issue of multimodal CL, we build on the feature covariance-aware metric (FeCAM) algorithm by Goswami et al. and propose an extended version of the algorithm that employs an incremental online semi-supervised learning process for each task. We call it exFeCAM or extended FeCAM algorithm. Additionally, our algorithm includes an intra-layer feature representation mechanism to learn more generalized feature maps for each class. To the best of our knowledge, we are the first to use a continual learning model in a multimodal data setting where an agent learns both tactile data and vision data provided by respective sensors. The main contributions of this article are summarized as follows:
  1. 1.
    We present an online non-exemplar continual learning algorithm with SSL capabilities. In addition, to enhance the feature maps from the pre-trained layers, we introduce intra-layer feature representations.
  2. 2.
    We present a new multimodal non-iid dataset for real world continual learning applications.
  3. 3.
    We demonstrate the robustness of learning different modalities as a new incremental domain rather than just combining the data from different modalities into a single fused feature vector.
  4. 4.
    We evaluate our algorithm on the custom multimodal dataset as well as the Core50 dataset Lomonaco and Maltoni and perform an ablation study to verify the effectiveness of the various parts of the algorithm.
  5. 5.
    We also perform a real-time experiment for object classification on the soft pneumatic gripper equipped with sensors and a separate camera setup, all synchronised using the ROS framework Koubaa et al.. More details about the experiment is provided in section 6.
The paper is structured as follows: In Section 2, an overview of the related works in CL, SSL, and multimodality is presented, with a focus on the recent advancements in the literature. Section 3 provides a detailed description of the experimental setup that has been adopted discussing the different cl strategies, data accumulation steps, and evaluation metrics. Following this, the proposed methodology is presented in Section 4, which includes the model architecture and the framework used in this study. The evaluation of the proposed methodology on custom multimodality dataset, as well as the Core50 dataset is provided in section 5. Moreover, in section 6 the paper showcases the real-time application of the proposed algorithm. Finally, a summary noting the advantages and disadvantages of the proposed method along with the scope for improvements and possible developments in the future are stated in the last section.
2 Related Work
2.1 Continual Learning Strategies
Continual learning strategies can be broadly categorized into four major types: architectural, regularization, rehearsal, and hybrid strategies, each with its own advantages and disadvantages. Architectural strategy either employs a dynamic modular network where new layers gets added upon encounter of new tasks Rusu et al., Fernando et al., Rajasegaran et al., Aljundi et al. or performs parameter isolation to allocate a dedicated separate sub-space for each task within the same network Mallya et al., Serra et al., Mallya and Lazebnik, Ahn et al.. The simplest form of architectural regularization involves freezing certain weights in the network to ensure they remain unchanged during training on new tasks. A major drawback of the architectural strategy is the increase in the size of the architecture as new tasks are learned or the potential saturation of the architecture if a fixed model is used.
Regularization strategies can be grouped into two: weight regularization and function regularization. Weight regularization constrains the update of the network weights by adding explicit regularization terms to the loss function, penalizing changes in each parameter based on its importance Kirkpatrick et al., Zenke et al., Chaudhry et al., Aljundi et al.. Function regularization, on the other hand, employs distillation methods to maintain the knowledge between the various layers of the network. This strategy typically employs the previously-learned model as the teacher and the currently model as the student, while implementing knowledge distillation (KD) techniques Hinton et al., Li and Hoiem, Rebuffi et al. Dhar et al., Lee et al.. However, regularising the model can make it more challenging for the network to acquire new knowledge.
Rehearsal Lopez-Paz and Ranzato, Bang et al., Kumari et al., Ebrahimi et al., Saha and Roy & pseudo-rehearsal, Shin et al., Wu et al., Ayub and Wagner, Pfülb and Gepperth, Wang et al. approaches offer the most simple yet effective approach by storing a subset of the previously learned data in a memory buffer or by training a generative model to account for the previous information and replaying it during training on new tasks. However, storing the images in a replay buffer for rehearsal can raise privacy concerns. On the other hand, while using a generative model somewhat addresses the privacy issues but may result in increased training time and computational demands.
Hybrid strategies Van de Ven et al., Liu et al., Ye and Bors are a combination of any of the other three strategies to compensate for catastrophic forgetting. SynapNet Kushawaha et al. uses a triple model architecture incorporating knowledge distillation along with a VAE-based pseudo-episodic memory for rehearsal and a sleep phase for memory re-organization. DualNet Pham et al. uses a fast learner for supervised learning and a slow learner for unsupervised learning of task-agnostic general representation using semi-supervised learning. ICARL Rebuffi et al. includes an external fixed memory to store a subset of old task data based on an elaborated sample selection procedure and then employs a distillation step acting as a regularization.
Non-exemplar continual learning strategies Goswami et al., Petit et al., Smith et al., Hayes et al. use the idea of replay-free sequential training, instead of using a modular network or regularizing the optimization function. Exemplar-free approaches store the prototypes of the past and new classes to compensate for catastrophic forgetting of the previously acquired knowledge. A pre-trained feature extractor provides the feature maps which are then used to calculate the mean and the covariance matrix for the respective classes.
2.2 Semi-Supervised Learning
Semi-supervised learning describes a class of algorithms designed to learn from both labeled and unlabeled data, typically assumed to be sampled from the same or similar distributions Yan et al., Mawuli et al.. One of the critical challenges in semi-supervised learning is how to effectively utilize labeled and unlabeled data while avoiding overfitting, preventing the effects of noise on the model, and ensuring effective model performance enhancement. Initial breakthroughs in semi-supervised learning with deep neural networks leveraged generative models such as autoencoders Rasmus et al., variational autoencoders Kingma et al., and generative adversarial networks Odena.
Some approaches in semi-supervised learning Oliver et al., Berthelot et al., Lee et al., Kingma et al., Kuo et al. also involve balancing a supervised loss ($l_{s}$) applied to the labeled data and an unsupervised consistency regularization loss $l_{ul}$ computed on the unlabeled data. Consistency regularization aims to improve model resilience and maintain label distribution, even in the presence of noisy images. It measures the discrepancy between predictions made on perturbed unlabeled data points. Approaches of this kind include the $\pi$-model Laine and Aila which augments the input data with image-based noise and self-regularizes through an additional consistency loss, mean teacher Tarvainen and Valpola, which employs an exponential moving average of parameters to regularize the model. Recently, fast-SWA Athiwaratkun et al. showed improved results by training with cyclic learning rates and measuring discrepancy with an ensemble of predictions from multiple checkpoints.
Two additional important approaches for semi-supervised learning, which have shown success both in the context of deep neural networks and other types of models are: self-training and conditional entropy minimization. Self-training also known as pseudo-labeling Lee et al. involves assigning classes to unlabeled data by making predictions from a model trained only on labeled data. Furthermore, co-training Xie et al. attempts to analyze data from different perspectives to solve the accumulative mistake issue during self-training. On the other hand, in case of conditional entropy minimization Grandvalet and Bengio, all unlabeled examples are encouraged to make confident predictions on some class. This doesn’t necessarily mean the predictions are correct (because the true labels for the unlabeled data are unknown), but the model is incentivized to be decisive rather than uncertain. In this paper, we use the concept of cosine similarity between labeled and unlabeled feature representations for pseudo-labeling of the unsupervised data. We refer to Van Engelen and Hoos for a comprehensive review on the topic.
2.3 Multimodality
Our experience of the world is inherently multimodal we perceive objects through vision, hear sounds, feel textures, smell odors, and taste flavors. To develop an AI agent with similar capabilities, researchers have explored various methods Donato et al., Babadian et al..
The simplest method to compensate multiple domains is using the concept of joint representation where the unimodal signals are combined in the same representation space by simple concatenation Baltrušaitis et al.. Ngiam et al. used autoencoders for multimodal domains. They employed stacked denoising autoencoders to represent each modality individually and subsequently fusing them into a unified multimodal representation using an additional autoencoder layer.
Similarly, Silberer and Lapata suggested a multimodal autoencoder for semantic concept grounding. They incorporated a reconstruction loss for training the representation and added a term in the loss function to predict object labels using the representation. Srivastava and Salakhutdinov introduced multimodal deep belief networks (DBNs) and multimodal deep boltzmann machines (DBMs) Srivastava and Salakhutdinov as multimodal representations. Kim et al. used a deep belief network for each modality and then combined them into joint representation for audiovisual emotion recognition. Huang and Kingsbury used a similar model for audio-visual speech recognition (AVSR), and Wu and Shao used it for audio and skeleton joint based gesture recognition. Ouyang et al. explored the use of multimodal DBMs for the task of human pose estimation from multi-view data. They demonstrated that integrating the data at a later stage after unimodal data, underwent nonlinear transformations, which was beneficial for the model. An alternative to a joint multimodal representation is a coordinated representation where instead of projecting the modalities together into a joint space, separate latent representations are learned for each modality and are coordinated through a constraint Frome et al..
Sarfraz et al. recently utilized a multimodal CL dataset to learn complementary information from two different modalities. They employed the VGGSound dataset, which includes 10-second video clips along with their corresponding audio. On the other hand, Xu et al. developed an egocentric multimodal dataset using smart glasses equipped with sensors, combining data from cameras, accelerometers, and gyroscopes. This dataset covers 32 types of daily activities performed by 10 participants. They utilized the temporal binding network (TBN) Kazakos et al. to fuse the information from different sensors into single feature vector for further training and feature extraction.
One of the major drawbacks of these methods is their inability to incrementally learn new domains. For example, if we train our model on two domains (sensor signals and images) and use a joint representation to learn these two modalities, introducing a new data modality later would necessitate re-training the model from scratch. Conversely, with our proposed exFeCAM model, we can add the new modality as a different domain and incrementally train the model without starting the training process anew. Additionally, our algorithm eliminates the necessity of storing data in memory for rehearsal, effectively addressing privacy concerns and storage challenges commonly encountered in exemplar-based CL methods.
With recent advancements in large language models (LLMs) and the introduction of GPT-4 Achiam et al. and its variants, it is now possible to utilize the data coming from vision, audio, and text signals into a single model. However, these models face significant challenges, including high computational and memory demands and a lack of continual learning capabilities.
Figure 1: Multimodal data comprising of two different domains, domain 1 contains the sensor information for different objects and domain 2 contains the images of those objects in different orientations, lightning conditions and different level of occlusion by the operator. Moving from left to right the model trains in a class-incremental manner, while going from top-bottom it trains in a domain-incremental manner.
3 Experimental Setup
In this section, we describe the steps involved in the creation of the multimodal continual learning dataset as well as the standard evaluation criteria for the class incremental as well as domain incremental learning scenarios.
3.1 Continual Learning Scenario
According to the division of incremental batches and the availability of task identities, continual learning can be broadly classified into three primary types or scenarios Van de Ven et al.:
  • Task-Incremental Learning (TIL): It requires an algorithm to learn a set of distinct tasks progressively. Task identities/numbers are provided in both training and testing phase.
  • Class-Incremental Learning (CIL): The algorithm must learn to differentiate between a growing number of classes incrementally. Task identities are only provided during training phase.
  • Domain-Incremental Learning (DIL): The algorithm needs to learn the same tasks but in different domains. Task identities are not required.
In addition to these major scenarios, there are also other scenarios present in the literature like task-free continual learning (TFCL), online continual learning (OCL), blurred boundary continual learning (BBCL), continual pre-training (CPT), we refer to Wang et al. for a thorough study.
Figure 2: Commercial two fingers pneumatic gripper equipped with two force sensors and two flex sensors affixed together using stretchable nylon thread. The flex sensor measures the bending of the finger and the force sensor measures the magnitude of the force applied to the finger tip.
3.2 Multimodal Dataset
The multimodal dataset comprises two components: sensor signals from an actuated finger and non-stationary images of objects captured by a vision camera. The two different modalities are fused into a hybrid setup that incorporates both class incremental as well as domain incremental scenarios as illustrated in Figure 1. In this setup, progression from left to right represents class-incremental training, while progression from top to bottom represents domain-incremental training. Both domains encompass five tasks, with each task involving the learning of two new classes. The first domain focuses on sensor signal data for various objects encountered in each experience and follows a class-incremental learning approach. The second domain deals with visual data of the same objects and employs both domain-incremental and class-incremental learning methods. Further details on data collection for the different modalities are provided in the subsequent sections.
3.2.1 Sensor Signal
To collect the sensor signals for the different objects we utilize a soft pneumatic gripper Hao et al. equipped with four sensors which are affixed together using stretchable nylon thread as illustrated in Figure 10. We have employed the commercially available flex sensors Saggio et al., Mishra et al. to measure the bending of the finger and a force sensitive resistor (FSR) Hollinger and Wanderley to quantify the force applied to the fingertip as shown in Figure 2. The outputs from the flex and force sensors are directed to an Arduino Due board, which in turn is connected to a computer for further processing and analysis. We gather the sensor data for 10 distinct objects, for each object the gripper acquires 50 data points, considering various orientations of the object. During the data collection process, for each data point the gripper holds and releases the object repeatedly at different contact points, with each cycle lasting $\sim$ 39 seconds. Each data point is represented as a 600-dimensional vector, encapsulating the sensor signals recorded during that time interval. The dataset acquisition for every object takes about 37 minutes to complete on a computer equipped with an Intel Core i7 CPU @2.80GHz (GTX 1070 GPU) running Windows 10.
3.2.2 Image Data
The image data consists of the same classes but repeated with three different objects. Each object is repeated 3 times to create a more diversified dataset. Classification can be performed at object level (30 classes) or at category level (10 classes). For each trial, the images are extracted from a 15-second video sequence at a rate of 10 frames per second (FPS). Objects are hand-held by the operator and the camera point-of-view is that of the operator’s eyes. The operator extends his arm and smoothly rotates the object in front of the camera. The grabbing hand (left or right) changes throughout the sessions and relevant object occlusions are often produced by the hand itself. This creates a complex non-stationary dataset with variable lighting conditions and arbitrary amounts of occlusion produced by the operator’s hand. The final dataset consists of 13,500 RGB images each of size 128x128 extracted from the video frame. The images are further resized to 32x32 dimension to speed up the training and evaluation process.
3.3 Evaluation Metrics
The performance of the CL algorithm depends on its ability to adapt, retain, and generalize the knowledge over time. We evaluate the incremental accuracy (denoted as $a_{t}$) of the model after every task as well as the average accuracy over all the tasks for each domain denoted as $A_{T}$. The average accuracy over all the domains is referred to as $A_{D}$ in the paper. We define the expressions as follows:
$A_{T}=\frac{1}{T}\sum_{t=1}^{t=T}a_{t}\\ A_{D}=\frac{1}{D}\sum_{T=1}^{T=D}A_{T}$ (1)
where $T$ and $D$ refer to the total number of incremental tasks in each domain and the total number of domains.
To evaluate the SSL capabilities of the algorithm we perform two different experiments as proposed in the paper Smith et al., we use the following terminology to describe them: First, Random Images, where we have 20% of unlabeled data and 80% labeled data for each class. To collect the unsupervised data we randomly sample 20% (about 240) data from the training dataset for each class, remove the class label, shuffle, and store it in a separate folder. Second, Unseen Images, where we sample the images of an unseen object of the same parent class from the training dataset and repeat the above steps. A point to note is that in the first case, the model might have already seen all the objects for each class, however for the second case, the model has never seen the objects present in the unsupervised dataset and thus makes it a more challenging situation. In addition, we also perform an ablation study to show the effectiveness of the intra-layer feature representation as well as SSL on the model’s overall accuracy.
Table 1: Pre-train model architecture for both the domains of the custom multimodal dataset, domain 2 represents the image data and domain 1 signifies the sensor data. The linear layers for both the models are removed once the pre-training is completed . Here, "classes" refers to the total number of output classes during the pre-training phase.
Model type Layer type Activation function Output channel/neurons Batch Normalization Other details
Conv2D Relu 16 Batchnorm2D kernel:3, stride:1, padding:1
Conv2D Relu 32 Batchnorm2D kernel:3, stride:2, padding:1
Conv2D Relu 64 Batchnorm2D kernel:3, stride:2, padding:1
Conv2D Relu 64 Batchnorm2D kernel:3, stride:3, padding:0
Domain 2 Conv2D Relu 128 Batchnorm2D kernel:2, stride:2, padding:0
Flatten - - - -
Linear Relu 2000 - -
Linear Relu 1000 - -
Linear Relu classes - -
Conv1D Leakyrelu (0.01) 512 Batchnorm1D kernel:3, stride:1, padding:1
Conv1D Leakyrelu (0.01) 256 Batchnorm1D kernel:3, stride:2, padding:1
Conv1D Leakyrelu (0.01) 128 Batchnorm1D kernel:3, stride:2, padding:1
Domain 1 Flatten - - - -
Linear Relu 256 - -
Linear - classes - -
4 Methodology
In this section, we discuss the modifications made to the FeCAM algorithm Goswami et al.. Originally, the algorithm models the feature distribution of classes using a multivariate normal distribution, $\mathcal{N}(\mu,\Sigma)$, and uses the Mahalanobis distance to perform classification. While the authors proposed several techniques to enhance and stabilize Mahalanobis-based distance classification, one significant limitation of FeCAM is its requirement to store batches in temporary memory and conduct training only once all data for a particular task is available. This makes it unsuitable for online learning. To address this issue and improve the algorithm’s capability for multimodal training, as well as to adapt to the semi-supervised nature of the dataset, we introduce several new features to the exFeCAM algorithm, which are detailed below.
4.1 Online Batchwise Training
Online training involves training a model incrementally in a batch-wise manner rather than on the entire dataset for a particular task at once. In this approach, each batch of data is passed through a feature extractor to obtain feature maps, which are then used to calculate the prototypes (mean and co-variance matrix) for each class in that batch. A dictionary is initialized with class labels as keys and their corresponding prototypes as values. There are two possible scenarios when updating this dictionary:
  1. 1.
    If a class is encountered for the first time, a new entry is added with the class label as the key and the prototypes as the value.
  2. 2.
    If the class has been seen before, the existing prototypes for that class are retrieved and updated by averaging the previous prototypes with the new ones from the current batch.
This method avoids the need to store incoming data for each task, which is critical for continual learning (CL) objectives and is especially suitable for autonomous agents with limited cache memory. Learning prototypes in a batch-wise manner addresses the challenge of not having access to previous data points, making it a more practical solution for real-time applications.
Algorithm 1 exFeCAM Algorithm
Input: Data stream $(D_{0}^{0},D_{0}^{1},...,D_{1}^{0},D_{1}^{1},...,D_{n}^{0:t})$, where "$t$" refers to the number of tasks, "$n$" refers to the number of domains.
Models: Feature extractor $f_{e}(x;\theta_{e})$, exFeCAM model $\phi(x;\mu,\Sigma)$
Variables: $k$ : number of layers of feature extractor, $t$: total tasks, $n$: number of domains, $T$: total classes; $b$: batch size; $l,j$ : iterative variables
$(x_{i},y_{i})\in D_{train}^{Supervised}$, $(x_{i}^{{}^{\prime}})\in D_{train}^{Unsupervised}$ , $(x_{i}^{{}^{\prime\prime}},y_{i}^{{}^{\prime\prime}})\in D_{test}$
[…]
Training Phase
for $E_{i}\in[E_{0},E_{1},...,E_{t}]$ do $\triangleright$ Different tasks/experiences
     Temporary Buffer $(\epsilon)\leftarrow\{\}$triangleright Temporary memory buffer for feature matching
     for $y_{i}\in[y_{0},y_{1},...,y_{T}]$ do $\triangleright$ Classes in each task
         for $j\leftarrow$ $b$ do
              $\psi_{ij}\leftarrow$ $[f_{e}(x_{i})_{k-(l-1)},f_{e}(x_{i})_{k-(l-2)},...,f_{e}(x_{i})_{k}]$triangleright Concatenate intra layer feature
              $\mu[i]\leftarrow\frac{1}{b}\sum\psi_{ij}$triangleright Mean matrix
              $\Sigma[i]\leftarrow$ Covariance($\psi_{ij}$) $\triangleright$ Covariance matrix
              $\Sigma[i]_{S}\leftarrow$ Shrinkage($\Sigma[i]$)
              $\Sigma[i]_{N}\leftarrow$ Normalization($\Sigma[i]_{S}$)
              if $\epsilon$ is not empty then
                  similarity $\leftarrow$ cosine similarity(features $\in\epsilon,x_{i}^{{}^{\prime}}$) $\triangleright$ Compute cosine similarity
                  if similarity $\geq$ threshold  then
                       Update $\mu[i]$, $\Sigma[i]_{N}$triangleright Update the mean and cov matrix for each class
                  end if
              end if
         end for
     end for
end for
Test Phase
for $x_{i}^{{}^{\prime\prime}}\in D_{test}$ do
     $\phi(f_{e}(x_{i}^{{}^{\prime\prime}});\mu,\Sigma)\leftarrow(f_{e}(x)-\mu_{i})^{T}$ $(\Sigma_{i})_{N}^{-1}$ $(f_{e}(x)-\mu_{i})$triangleright Calculate the Mahalanobis distance
     $y_{pred}\leftarrow$ argmin $\phi(f_{e}(x_{i}^{{}^{\prime\prime}});\mu,\Sigma)$ $\forall$ $y_{i}=0,1,...T$triangleright Assign the class with minimum Mahalanobis distance
end for
4.2 Intra Layer Feature Representation
Intra-layer feature representation (ILFR) leverages the power of intermediate layers that encodes the low and mid-level feature information present in the data extracted using a pre-trained model. Instead of relying solely on the final output layer of the feature extractor, the presence of multiple activation maps generalizes the learned knowledge and further improves the performance of the algorithm. We argue that enriching the last layer representations with hierarchical intra-layer features increases robustness to domain shifts and thus improves generalizability to downstream continual tasks. These extracted features can be the class embeddings of a transformer encoder Vaswani et al. or flattened feature maps of a ResNet encoder He et al..
There are two intuitive ways to promote intra-layer feature representations into the model: averaging the representations from the last "k" layers of the network or concatenating the representations from the last k layers. The first approach requires the output dimensions of all layers to be identical to facilitate summation. In our use case, we opt for concatenation, as the outputs from different layers have varying dimensions. Concatenation also preserves the rich, multi-scale features, unlike summation, which can result in information loss due to the merging of complementary features captured by different layers, potentially reducing the richness of the learned representations.
For the multimodal experiment, we pre-train the domain 1 feature extractor on gripper data of objects not present in the training dataset, while the domain 2 feature extractor is trained on the Core50 dataset. More details about the pre-trained model architecture are provided in Table 1. For the Core50 experiment, we use the ResNet18 architecture as the feature extractor, pre-trained on the ImageNet1k dataset Russakovsky et al..
4.3 Semi-Supervised Learning
Semi-supervised learning (SSL) denotes the capability of an algorithm to learn from supervised data and then improve the performance of the model by training it on the unsupervised part. It is very useful in a situation where we have limited supervised data but we can collect huge amounts of unsupervised data from the environment. This is an active area of research given that large, labeled datasets are expensive, but most applications have access to plentiful, cheap unlabeled data. In this paper, we use the concept of pseudo-labeling to organize the unsupervised data and provide them with dummy labels for further training. To perform pseudo labeling we randomly store the feature maps of each class for every task in a temporary memory buffer and then perform cosine similarity between the reference feature maps and the feature maps generated from the complete unsupervised dataset as defined in equation 2,
$similarity=\frac{x_{1}.x_{2}}{||x_{1}||_{2}||x_{2}||_{2}}$ (2)
where $x_{1}$ and $x_{2}$ refers to the reference feature map present in the temporary memory and the feature maps from the unsupervised dataset, ||.$||.||_{2}$ refers to the l2 norm of the two tensors. If the value of the cosine similarity is greater than a certain threshold, we then train the exFeCAM algorithm on the respective feature maps and its associated pseudo labels. These pseudo-labels are derived from the class labels used for feature matching. The prototype values for that class are subsequently updated using the averaging rule described in the previous section. To determine the threshold value, we employ the Optuna framework Akiba et al., which uses bayesian optimization to explore the hyperparameter space efficiently.
Figure 3: Modified FeCAM algorithm with intra layer feature representation, online batchwise learning, and semi-supervised learning capabilities for multimodal training. During training on multiple domains, the algorithm stores the mean and covariance matrices of the classes in a dictionary. In the testing phase, the algorithm uses these stored prototypes and employs the Mahalanobis distance to determine the prediction class. Supervised training is depicted with solid black lines, while unsupervised training is illustrated with dashed black lines.
4.4 Multimodal Training
We extend the class-incremental continual learning to a more realistic semi-supervised continual learning (SSCL) setting, where data distributions reflect existing object class correlations between, and among, the labeled and unlabeled data distributions. For our experiment we use 70% labeled data and 30% unlabeled data for the two protocols (random and unseen) discussed above. At task n, we denote batches of labeled data as $X_{n}={(x_{b},y_{b}):b_{i}\in(1,...,b)|y_{b}\in\tau_{n}}$ and batches of unlabeled data as $X_{n}^{{}^{\prime}}={(x_{b}^{{}^{\prime}}):b_{i}\in(1,...,b)}$, where $b$ refers to the batch size. The goal in task $n$ is to learn a model which predicts object class labels for any query input over all classes seen in the current and previous tasks ($\tau_{0}\cup\tau_{1}\cup...\cup\tau_{n}$). During each training step the model receives different classes incrementally over all the experiences for each domain as shown in Figure 1. The two domains contain data from different modalities specifically domain 1 contains the sensor data whereas domain 2 contains the vision data. The data for each task goes through a feature extractor to derive the feature maps from the different layers of the network. The feature maps for both domains are standardized to a fixed dimension of 384 and normalized to fall within the 0-1 range before being fed into the subsequent stages of the algorithm. These feature maps are then used to incrementally train the exFecam algorithm. During training for each task in the second domain, a small subset of the feature maps of the object is stored in a temporary memory buffer to facilitate pseudo-labeling for semi-supervised learning, as previously discussed.
The feature maps are used to compute the mean and covariance matrix for each batch during training. The model performs covariance matrix approximation, shrinkage, and normalization to stabilize the training process Goswami et al.. During the test phase, the extracted feature maps from the test dataset are used to calculate the Mahalanobis distance for all seen classes, and the class with the minimum distance value is selected as the output class. More details about the implementation of the algorithm is provided in the pseudo code. Additionally, we also perform a real-time test for object classification based on the multimodal data using the ROS framework as discussed in section 6 of the paper.
Figure 4: Intra layer feature representation for the custom dataset and the Core50 dataset, we compare the improvement in the accuracy of the FeCAM algorithm on addition of knowledge from the last as well the penultimate layers of the pre-trained feature extractors. The improvement in case of Core50 dataset is considerably more than the custom dataset due to the presence of more diverse information in case of image data.
We further analyzed the intra-layer feature representation for both the custom dataset and the Core50 dataset by comparing the accuracy improvement of the exFeCAM algorithm when incorporating knowledge from the last and penultimate layers of pre-trained feature extractors. The Core50 dataset showed a significantly greater improvement in accuracy compared to the custom dataset. The difference is attributed to the more diverse information contained in the image data of the Core50 dataset.
Figure 5: On the left: radar plot shows the accuracy of the two domains after incremental training on different tasks. The FeCAM algorithm learns the first domain in a class-incremental approach, while it learns the second domain in a combined domain-incremental and class-incremental manner. On the right: the plot shows domain 1 accuracy (in orange) with random gaussian noise added to it, the solid line in grey shows the percentage improvement in the overall accuracy of the experiences after addition of the second domain. The black lines show the error bars for the two cases.
5 Results
To assess the performance of our algorithm and the effectiveness of the proposed modifications, we evaluate the exFeCAM algorithm on a custom multimodal dataset and the standard Core50 dataset used in the literature. The Core50 dataset was chosen for its non-iid nature and its compatibility for self-supervised learning. We do not explicitly compare our algorithm with other state-of-the-art (SOTA) benchmarks because the focus of this paper is not to establish a new level of accuracy for object classification, but rather to demonstrate the effectiveness of CL in a real world multimodal training setting. Nevertheless, our results are based on improvements to the original FeCAM algorithm, indirectly providing a comparison. All experiments are repeated three times, and we plot the mean and standard deviation of the respective metrics.
Figure 6: Accuracy of the FeCAM algorithm on SSL tasks, the blue and green plot shows the SSL condition for unseen object from the same parent class, with blue indicating SSL enabled and green indicating SSL disabled case. The maroon and red plots represent the SSL condition for randomly selected images of the object, with maroon indicating SSL enabled and red indicating SSL disabled case. Top plot shows the accuracy for the custom multimodal dataset and the botton plot shows the accuracy for the Core50 dataset. The black dotted lines denotes the average accuarcy over all the experiences for different SSL cases.
[…]
Figure 7: The real-time multimodal experiment using ROS involves three publishers and one subscriber node. The publishers (P1, P2, P3) are depicted in blue, while the subscriber is shown in yellow. P1 operates at a frequency of 20 Hz, P2 operates at 10 Hz, and P3 operates at 0.2 Hz. The output from P3 is used to actuate a soft finger equipped with sensors. The outputs from P1 and P2 are fed to the subscriber, where they undergo pre-processing, feature extraction using a pre-trained network, and finally to the FeCAM algorithm to make the predictions using the saved mean and co-variance matrices.
Figure 4 (bottom) presents similar information for the Core50 dataset, where we employed the same experimental protocols. The key difference lies in the pre-trained model and the pre-training dataset used in this experiment. However, the improvement in accuracy for the Core50 dataset is minimal compared to the custom dataset. This could be attributed to the significant variations in the objects within each class, making the simple cosine similarity method less effective in this context. A more complex SSL method might be more beneficial for this scenario, which we plan to explore in future experiments.
6 Real-Time Evaluation With ROS
Robot Operating System (ROS) is a comprehensive framework that comprises a suite of tools, libraries, and protocols aimed at facilitating the development of various robotic systems. The system manages the creation and control of communication between a robot’s peripheral modules, such as sensors, cameras, and actuators, thereby enabling the seamless integration of these components with parallization capabilties.
Figure 8: Real-time images of the object from the camera’s perspective, along with the sensor data from four sensors attached to the soft finger, upon its actuation. These images and sensor readings are directly fed into the FeCAM model after resizing them to the desired dimension. For a single 15-second run, there are a total of 150 sensor signals and 300 images, representing the number of raw data points and corresponding images of the object.
6.1 ROS Setup
ROS is primarily designed to function optimally on Ubuntu, while only partially compatible with other operating systems, such as Windows and Mac. For our case we use ROS on Ubuntu 20.04 on a Intel Core i7 7th gen CPU @@@@ 2.80 GHz equipped with Nvidia GTX 1080 GPU. We use the Noetic Ninjemys distribution of ROS 1 Koubaa et al. in our experiment. A detailed explanation about the workflow of the project is provided in the flowchart in Figure 7.
Figure 9: Square wave sent to the control box to actuate the soft pneumatic finger. We pressurise the finger for 3 seconds with a maximum pressure of 1.3 bars and then depressurise it for 2 seconds, the cycle is repeated "N" times till we reach the 15 seconds threshold for each run.
The first publisher (P1) publishes the sensor signals coming from the four sensors attached to the soft finger, the second publisher (P2) publishes the video coming from the camera and the third publisher (P3) publishes the command pressure to the control box which inturn actuates the finger. The control frequency of P1 and P2 are 10hz and 20hz, whereas the actuation frequency of P3 is 0.2hz. We further observed that actuating the finger at a more higher frequency constrained it from properly holding the object and thus effected the quality of the generated data from the sensors. The data from P1 and P2 goes to the subscriber which then pre-process the data, extracts the frames from the video in case of P2 and sends it to the respective feature extractors. The output of the feature extractor is a 384-dimensional feature map (in case the intra-layer feature representation is turned on) which goes as an input to the exFecam model for further processing.
6.2 Real-Time Experiment
To assess the performance of our algorithm, we conduct a quasi real-time experiment using a soft pneumatic gripper equipped with two flex sensors and two force sensors. During the evaluation phase, data from the sensors and the camera are collected in real-time for approximately 15 seconds. These data are treated as two distinct domains and are sent through the pipeline in parallel for further processing and subsequently to the exFecam model for predictions. To generate the predictions we utilize the stored mean and the co-variance matrix from a prior offline incremental experiment conducted for the two domains. Impressively, the mean and covariance matrix occupy only 22.5 MB of memory, making it quite efficient compared to other continual learning strategies
However, since the model was initially trained on images of the objects in a different frame of reference, where the operator extends his arm and smoothly rotates the object in front of the camera. In contrast, during the real-time experiment, the objects are positioned near the soft finger, and a camera captures images from a top-down view. This difference sometimes leads to incorrect predictions by the model. To address this issue, for some objects, we occasionally need to refine the model’s knowledge by conducting a brief training session (lasting around 82 seconds) specifically for that object. Figure 8 shows a snapshot of the real-time data collection from both the sensors and the camera’s perspective. The image shows the progression of the sensor signal and the camera image upon actuating the finger. For each cycle the sensors collect 150 raw values, which are transformed into a single 600-dimensional vector, whereas the camera collects around 300 images of the object from various orientations within the 15-seconds interval.
As discussed above one of the major advantages of this domain incremental training is for autonomous agents in situations where the field of view of one of the sensors gets restricted due to the presence of some obstacle from the environment, in that case the presence of other sensor can compensate for the missing knowledge.
6.3 Real World Application
A real world application for such a technique can be in companion robots Broadbent et al., where the robot learns new tasks incrementally by interacting with its environment and recognizing different objects using information from multiple sensors available to it. The use of multiple sensors is particularly advantageous in scenarios where a single sensor cannot provide complete information about an object. For instance, a vision camera might struggle to differentiate between two objects with similar structures but with different textures or materials, with one being much softer than the other. In such cases, information from a tactile sensor can be extremely beneficial. Our proposed framework simplifies the addition of new sensors by treating each as a new domain, eliminating the need to retrain the entire algorithm from scratch each time. All the sensors and the algorithm can be synchronized using a mobile version of ROS for embedded systems Min et al., providing a comprehensive solution.
7 Conclusion and Discussion
In this paper, we present a novel continual learning method that combines the informations from different sensors in a domain incremental fashion while also learning new classes or objects in a class incremental manner. We enhanced the original FeCAM algorithm by incorporating online batch-wise training, intra-layer feature representation, and SSL capabilities thereby making the algorithm suitable for our scenario. The exFeCAM algorithm employs cosine similarity based feature matching to perform pseudo labeling when encountering unlabeled data. We validated the algorithm on two different modalities of data namely sensor signal based data and image based data, and demonstrated the ability of the algorithm to learn the two modalities incrementally. We also conducted an ablation study to show the contribution of different parts of the algorithm. Additionally, we performed a real-time experiment on a soft pneumatic gripper equipped with four sensors and an external camera system, all synchronised using the ROS framework. As part of our future work, we aim to enhance the exFeCAM algorithm further by adding forward and backward knowledge capabilities, either through interactions between class prototypes or by training a simple multilayer perceptron (MLP) network to learn these interactions. Moreover, currently we need to manually select the prediction domain modality in the presence of noise; in the future, we aim to enable the algorithm to automatically determine this by calculating the confidence of each prediction from the network.
CRediT authorship contribution statement
Nilay Kushawaha: Writing - original draft, Methodology, Investigation, Experimentation, Formal Analysis, Conceptualization. Egidio Falotico: Supervision, Funding acquisition, Project administration.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request. The code is available on the github repo : https://github.com/nilay121/Multimodal-Fusion-using-CL
Acknowledgment
We acknowledge the contribution from the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union–NextGenerationEU (Project IR0000011, CUP B51E22000150006, "EBRAINS-Italy")

### v2

@@
-Conceptualization of this study, Methodology, Software
+Continual Learning for Multimodal Data Fusion of a Soft Gripper
 Abstract
-Continual learning (CL) refers to the ability of an algorithm to continuously and incrementally acquire new knowledge from its environment while retaining previously learned information.
-A model trained on one data modality often fails when tested with a different modality.
-A straightforward approach might be to fuse the two modalities by concatenating their features and training the model on the fused data.
-However, this requires retraining the model from scratch each time it encounters a new domain.
-In this paper, we introduce a continual learning algorithm capable of incrementally learning different data modalities by leveraging both class-incremental and domain-incremental learning scenarios in an artificial environment where labeled data is scarce, yet non-iid (independent and identical distribution) unlabeled data from the environment is plentiful.
-The proposed algorithm is efficient and only requires storing prototypes for each class.
-We evaluate the algorithm’s effectiveness on a challenging custom multimodal dataset comprising of tactile data from a soft pneumatic gripper, and visual data from non-stationary images of objects extracted from video sequences.
-Additionally, we conduct an ablation study on the custom dataset and the Core50 dataset to highlight the contributions of different components of the algorithm.
-To further demonstrate the robustness of the algorithm, we perform a real-time experiment for object classification using the soft gripper and an external independent camera setup, all synchronized with the Robot Operating System (ROS) framework.
+Continual learning (CL) refers to the ability of an algorithm to continuously and incrementally acquire new knowledge from its environment while preserving previously learned information.
+Traditional models trained on a single data modality often fail to generalize when exposed to different modalities.
+A simple strategy to address this limitation is to fuse multiple modalities by concatenating their features and training the model on the combined representation.
+However, this generally requires re-training the model from scratch whenever a new domain is introduced.
+In this paper, we propose a continual learning algorithm that incrementally learns from multiple data modalities by combining class-incremental and domain-incremental learning settings.
+We validate our algorithm on a custom multimodal dataset composed of tactile signals collected from a soft sensorized pneumatic gripper and visual data consisting of non-stationary object images captured from video sequences.
+Additionally, we evaluate our method on a subset of the publicly available VGGSound dataset, which integrates both visual and audio signals corresponding to various real-world activities.
+To further demonstrate the robustness and real-time applicability of our approach, we conduct an object classification experiment using a soft sensorized gripper and an external camera setup, all synchronized through the Robot Operating System (ROS) framework.
 keywords:
-[…]
+Continual Learning, Class prototypes, Multimodality, Robot operating system, Semi-supervised learning, Soft gripper, Tactile data
+Nilay Kushawaha* Egidio Falotico
+{affiliations}
+Nilay Kushawaha
+Address[1]: The BioRobotics Institute, Scuola Superiore Sant’Anna, Pontedera 56025, Italy
+Address[2]: Department of Excellence in Robotics and AI, Scuola Superiore Sant’Anna, Pisa 56127, Italy Email Address: nilay.kushawaha@santannapisa.it
+Egidio Falotico
+Address[1]: The BioRobotics Institute, Scuola Superiore Sant’Anna, Pontedera 56025, Italy
+Address[2]: Department of Excellence in Robotics and AI, Scuola Superiore Sant’Anna, Pisa 56127, Italy
 1 Introduction
-In recent years, machine learning models have matched or even exceeded human-level performance on tasks such as image classification Ahmad et al., object recognition Zou et al., natural language processing (NLP) Otter et al., playing games in simulation, etc.
-While these models excel in static tasks where the data distribution remains constant, they encounter difficulties in adapting to changing environments and need to restart the training process whenever new data is introduced.
-On the other hand, humans are adept at learning from dynamic environments, continuously acquiring, updating, and applying knowledge over time.
-We expect AI systems to adapt in a similar fashion.
-Continual learning (CL) Wang et al., Lesort et al., also known as lifelong learning or incremental learning, is a machine learning subfield that focuses on progressively training models on a continuous stream of data to accumulate and retain knowledge over time.
-However, a significant obstacle for these artificial models is their tendency to forget previously acquired skills when exposed to new tasks or data distributions.
-This issue, known as catastrophic forgetting, often leads to a substantial drop in performance as the new information overwrites the old knowledge or alters the data distribution.
-A naive solution to this problem is to save all data, shuffle it, and retrain the model offline from scratch.
-While this method effectively addresses catastrophic forgetting, it is time-inefficient, suboptimal, and unsustainable for modern large-scale models that require significant computational power.
-Continual learning aims to develop methods that balance performance and efficiency more effectively.
-Robots operating in unstructured real world environments will encounter new tasks and challenges over time, necessitating capabilities that cannot be fully anticipated from the outset.
-These robots need to learn continually, acquiring new skills without forgetting previously learned ones.
-A model trained in one domain will be ineffective when deployed in a completely different domain or used with a different data modality than it was originally trained on.
-One straightforward approach might be to use separate models for each data modality or to jointly train a single model on all modalities offline.
-However, this is not a feasible solution for agents with limited memory and computational resources that learns new domains through active exploration and interaction with its environment.
-Another challenge is the limited availability of labeled data for supervised incremental training.
-A large research effort is dedicated towards learning from unlabeled data, which is often more readily available than labeled data in practical applications.
-Within this effort, the field of semi-supervised learning (SSL) Van Engelen and Hoos has recently demonstrated the most promising results.
-In this setting, an agent can utilize a significant amount of unlabeled data from its environment.
-It can employ the knowledge gained from labeled data to learn new objects of the same class or to enrich its understanding of previously learned classes using the unlabeled data.
-In this paper, we consider a realistic continual learning problem where a single continual learning model learns different modalities of data incrementally by leveraging both class-incremental and domain-incremental learning scenarios in an artificial environment where labeled data is scarce yet non-iid (independent and identical distribution) unlabeled data from the environment is plentiful.
-We mainly experiment with two different SSL conditions: Radom images, where for each object of the same parent class we randomly sample some images from the dataset that are not included in the training data; Unique objects, where we sample the images of new unseen objects that belong to the same parent class.
+In recent years, machine learning models have matched or even exceeded human-level performance on various tasks, such as image classification [1], object recognition [2], natural language processing (NLP) [3], and simulated game play [4].
+Despite their success in static settings where data distributions remain stable, these models struggle to adapt to evolving environments and often require re-training the network from scratch when new data becomes available.
+In contrast, humans demonstrate remarkable adaptability, continually learning, updating, and applying knowledge over time.
+Our goal is to develop AI systems that can replicate this level of continual adaptability.
+Continual learning (CL) [5], also referred to as lifelong or incremental learning is a subfield of machine learning that focuses on training models progressively on a stream of data, allowing it to accumulate and retain knowledge over time.
+However, a major challenge for such models is their susceptibility to catastrophic forgetting of the previously acquired skills when learning new tasks/experiences or encountering shifts in data distribution.
+This often results in significant performance degradation, as new information overwrites existing knowledge.
+A straightforward but impractical solution is to store all past data, shuffle it, and re-train the model offline from scratch.
+While this can mitigate forgetting, it is computationally expensive, inefficient, and unsustainable for large-scale models.
+The goal of continual learning is to develop more scalable and efficient approaches that can maintain performance over time.
+A model trained in one domain often performs poorly when deployed in a completely different domain or when applied to a different data modality than it was originally designed for.
+A common approach to address this issue is to either train separate models for each modality or to jointly train a single model offline on all modalities.
+However, these strategies are impractical for agents with limited memory and computational resources, especially those that learn new domains through active exploration and interaction with their environment.
+Another significant challenge is the scarcity of labeled data for supervised incremental learning.
+In contrast, unlabeled data is typically more abundantly available in real-world applications.
+To address this imbalance, semi-supervised learning (SSL) [6] has emerged as a promising solution.
+In this setting, an agent leverages a small set of labeled data to extract useful knowledge, which can then be applied to learn new instances of the same class or to enhance it’s understanding of previously encountered classes using the unlabeled data.
+In this paper, we consider a realistic CL problem where a single CL model learns different modalities of data incrementally by leveraging both class-incremental and domain-incremental learning scenarios in an artificial environment where labeled data is scarce yet non-iid (independent and identical distribution) unlabeled data from the environment is plentiful.
+We also experiment with two different SSL conditions: Radom images, where for each object of the same parent class we randomly sample some images from the dataset that are not included in the training data; Unique objects, where we sample the images of new unseen objects that belong to the same parent class.
 More details about the SSL conditions is provided in section 3.3.
-To perform a multimodal CL experiment, we leverage two different modalities of data mainly the tactile data coming from sensors integrated with a soft gripper and the visual data from an independent camera setup.
-The CL model is designed to learn the classes in both the domains incrementally.
-To address this issue of multimodal CL, we build on the feature covariance-aware metric (FeCAM) algorithm by Goswami et al.
-and propose an extended version of the algorithm that employs an incremental online semi-supervised learning process for each task.
-We call it exFeCAM or extended FeCAM algorithm.
-Additionally, our algorithm includes an intra-layer feature representation mechanism to learn more generalized feature maps for each class.
-To the best of our knowledge, we are the first to use a continual learning model in a multimodal data setting where an agent learns both tactile data and vision data provided by respective sensors.
-The main contributions of this article are summarized as follows:
+For our multimodal CL experiment, we employ two distinct types of dataset.
+The first setup involves tactile input from sensors embedded within a soft gripper, alongside visual data captured by an independent camera system.
+For the second experiment, we use the VGGSound dataset [7], which offers paired image and audio modalities extracted from video clips.
+Our CL model is designed to incrementally learn class representations across these different domains.
+To address the challenges associated with multimodal CL, we build upon the Feature Covariance-Aware Metric (FeCAM) algorithm introduced by [8].
+We propose an enhanced version, referred to as exFeCAM (extended FeCAM), which incorporates an online, task-wise semi-supervised learning strategy.
+Furthermore, our approach introduces an intra-layer feature representation mechanism that enables the learning of more generalized feature maps for each class.
+To the best of our knowledge, this is the first work to apply continual learning in a multimodal setting, enabling an agent to learn from different modalities in a domain-incremental manner.
+The key contributions of this work are as follows:
   1. 1.
-    We present an online non-exemplar continual learning algorithm with SSL capabilities.
-In addition, to enhance the feature maps from the pre-trained layers, we introduce intra-layer feature representations.
+    We present an online non-exemplar CL algorithm with SSL capabilities. In addition, we incorporate intra-layer feature representation technique to obtain generalized feature maps.
   2. 2.
-    We present a new multimodal non-iid dataset for real world continual learning applications.
+    We present a new multimodal non-iid dataset for real-world CL applications.
   3. 3.
@@
   4. 4.
-    We evaluate our algorithm on the custom multimodal dataset as well as the Core50 dataset Lomonaco and Maltoni and perform an ablation study to verify the effectiveness of the various parts of the algorithm.
+    We evaluate our algorithm on the custom multimodal dataset as well as the VGGSound dataset [7] and perform an ablation study to verify the effectiveness of the various parts of the algorithm.
   5. 5.
-    We also perform a real-time experiment for object classification on the soft pneumatic gripper equipped with sensors and a separate camera setup, all synchronised using the ROS framework Koubaa et al..
+    We also perform a real-time experiment for object classification using a soft pneumatic gripper equipped with sensors and a separate camera setup, all synchronized using the ROS framework [9].
 More details about the experiment is provided in section 6.
@@
 Following this, the proposed methodology is presented in Section 4, which includes the model architecture and the framework used in this study.
-The evaluation of the proposed methodology on custom multimodality dataset, as well as the Core50 dataset is provided in section 5.
+The evaluation of the proposed methodology on the custom multimodality dataset, as well as the VGGSound dataset is provided in section 5.
 Moreover, in section 6 the paper showcases the real-time application of the proposed algorithm.
@@
 2.1 Continual Learning Strategies
-Continual learning strategies can be broadly categorized into four major types: architectural, regularization, rehearsal, and hybrid strategies, each with its own advantages and disadvantages.
-Architectural strategy either employs a dynamic modular network where new layers gets added upon encounter of new tasks Rusu et al., Fernando et al., Rajasegaran et al., Aljundi et al.
-or performs parameter isolation to allocate a dedicated separate sub-space for each task within the same network Mallya et al., Serra et al., Mallya and Lazebnik, Ahn et al..
-The simplest form of architectural regularization involves freezing certain weights in the network to ensure they remain unchanged during training on new tasks.
-A major drawback of the architectural strategy is the increase in the size of the architecture as new tasks are learned or the potential saturation of the architecture if a fixed model is used.
+Continual learning strategies can be broadly categorized into four major types: architectural, regularization, rehearsal, and hybrid strategies, each with it’s own pros and cons.
+Architectural strategy either employs a dynamic modular network where new layers gets added upon encounter of new tasks [10], [11] or performs parameter isolation to allocate a dedicated separate sub-space for each task within the same network [12], [13].
+However, a major drawback of this strategy is the increase in the size of the architecture as new tasks are learned or the potential saturation of the architecture if a fixed model is used.
 Regularization strategies can be grouped into two: weight regularization and function regularization.
-Weight regularization constrains the update of the network weights by adding explicit regularization terms to the loss function, penalizing changes in each parameter based on its importance Kirkpatrick et al., Zenke et al., Chaudhry et al., Aljundi et al..
+Weight regularization constrains the update of the network weights by adding explicit regularization terms to the loss function, penalizing changes in each parameter based on it’s importance [14], [15].
 Function regularization, on the other hand, employs distillation methods to maintain the knowledge between the various layers of the network.
-This strategy typically employs the previously-learned model as the teacher and the currently model as the student, while implementing knowledge distillation (KD) techniques Hinton et al., Li and Hoiem, Rebuffi et al.
-Dhar et al., Lee et al..
-However, regularising the model can make it more challenging for the network to acquire new knowledge.
-Rehearsal Lopez-Paz and Ranzato, Bang et al., Kumari et al., Ebrahimi et al., Saha and Roy & pseudo-rehearsal, Shin et al., Wu et al., Ayub and Wagner, Pfülb and Gepperth, Wang et al.
-approaches offer the most simple yet effective approach by storing a subset of the previously learned data in a memory buffer or by training a generative model to account for the previous information and replaying it during training on new tasks.
+This typically involves treating the previously trained model as a teacher and the current model as a student, applying knowledge distillation (KD) techniques to align their outputs [16], [17].
+However, regularizing the model can make it more difficult for the network to acquire new knowledge.
+Rehearsal [18], [19], & pseudo-rehearsal, [20], [21] approaches offer the most simple yet effective approach by storing a subset of the previously learned data in a memory buffer or by training a generative model to account for the previous information and replaying it during training on new tasks.
 However, storing the images in a replay buffer for rehearsal can raise privacy concerns.
-On the other hand, while using a generative model somewhat addresses the privacy issues but may result in increased training time and computational demands.
-Hybrid strategies Van de Ven et al., Liu et al., Ye and Bors are a combination of any of the other three strategies to compensate for catastrophic forgetting.
-SynapNet Kushawaha et al.
-uses a triple model architecture incorporating knowledge distillation along with a VAE-based pseudo-episodic memory for rehearsal and a sleep phase for memory re-organization.
-DualNet Pham et al.
-uses a fast learner for supervised learning and a slow learner for unsupervised learning of task-agnostic general representation using semi-supervised learning.
-ICARL Rebuffi et al.
-includes an external fixed memory to store a subset of old task data based on an elaborated sample selection procedure and then employs a distillation step acting as a regularization.
-Non-exemplar continual learning strategies Goswami et al., Petit et al., Smith et al., Hayes et al.
-use the idea of replay-free sequential training, instead of using a modular network or regularizing the optimization function.
-Exemplar-free approaches store the prototypes of the past and new classes to compensate for catastrophic forgetting of the previously acquired knowledge.
+While using a generative model somewhat addresses the privacy issues, it may result in increased training time and computational demands.
+Hybrid strategies [22], [23] are a combination of any of the other three strategies to compensate for catastrophic forgetting.
+SynapNet [24] uses a triple model architecture along with a VAE-based pseudo-episodic memory for rehearsal and a sleep phase for memory re-organization.
+DualNet [25] uses a fast learner for supervised learning and a slow learner for unsupervised learning of task-agnostic general representation using semi-supervised learning.
+Non-exemplar CL strategies [8], [26] use the idea of replay-free sequential training, instead of using a modular network or regularizing the optimization function.
+These approaches store the prototypes of the past and new classes to compensate for catastrophic forgetting of the previously acquired knowledge.
 A pre-trained feature extractor provides the feature maps which are then used to calculate the mean and the covariance matrix for the respective classes.
 2.2 Semi-Supervised Learning
-Semi-supervised learning describes a class of algorithms designed to learn from both labeled and unlabeled data, typically assumed to be sampled from the same or similar distributions Yan et al., Mawuli et al..
-One of the critical challenges in semi-supervised learning is how to effectively utilize labeled and unlabeled data while avoiding overfitting, preventing the effects of noise on the model, and ensuring effective model performance enhancement.
-Initial breakthroughs in semi-supervised learning with deep neural networks leveraged generative models such as autoencoders Rasmus et al., variational autoencoders Kingma et al., and generative adversarial networks Odena.
-Some approaches in semi-supervised learning Oliver et al., Berthelot et al., Lee et al., Kingma et al., Kuo et al.
-also involve balancing a supervised loss ($l_{s}$) applied to the labeled data and an unsupervised consistency regularization loss $l_{ul}$ computed on the unlabeled data.
-Consistency regularization aims to improve model resilience and maintain label distribution, even in the presence of noisy images.
-It measures the discrepancy between predictions made on perturbed unlabeled data points.
-Approaches of this kind include the $\pi$-model Laine and Aila which augments the input data with image-based noise and self-regularizes through an additional consistency loss, mean teacher Tarvainen and Valpola, which employs an exponential moving average of parameters to regularize the model.
-Recently, fast-SWA Athiwaratkun et al.
-showed improved results by training with cyclic learning rates and measuring discrepancy with an ensemble of predictions from multiple checkpoints.
-Two additional important approaches for semi-supervised learning, which have shown success both in the context of deep neural networks and other types of models are: self-training and conditional entropy minimization.
-Self-training also known as pseudo-labeling Lee et al.
-involves assigning classes to unlabeled data by making predictions from a model trained only on labeled data.
-Furthermore, co-training Xie et al.
-attempts to analyze data from different perspectives to solve the accumulative mistake issue during self-training.
-On the other hand, in case of conditional entropy minimization Grandvalet and Bengio, all unlabeled examples are encouraged to make confident predictions on some class.
-This doesn’t necessarily mean the predictions are correct (because the true labels for the unlabeled data are unknown), but the model is incentivized to be decisive rather than uncertain.
-In this paper, we use the concept of cosine similarity between labeled and unlabeled feature representations for pseudo-labeling of the unsupervised data.
-We refer to Van Engelen and Hoos for a comprehensive review on the topic.
+SSL describes a class of algorithms that aim to learn from a combination of labeled and unlabeled data, which are typically assumed to be drawn from the same underlying distributions [27].
+A central challenge in SSL lies in effectively leveraging both labeled and unlabeled data while mitigating the risks of overfitting, reducing the impact of noisy data, and ensuring meaningful improvements in model performance.
+Early advances in SSL using deep neural networks were driven by generative models such as autoencoders [28], variational autoencoders [29], and generative adversarial networks [30], which played a crucial role in extracting robust representations from partially labeled datasets.
+Some SSL approaches [31], [32] focus on balancing two complementary loss terms: a supervised loss ($l_{s}$) applied to labeled samples and an unsupervised consistency regularization loss ($l_{ul}$) computed on the unlabeled data.
+These approaches are designed to ensure that the model learns not only from explicit supervision but also from consistency in predictions across augmented versions of unlabeled inputs.
+Two additional important approaches which have shown success in the context of deep learning are: self-training and conditional entropy minimization.
+Self-training, also known as pseudo-labeling [33], involves assigning labels to unlabeled data based on the predictions of a model initially trained on labeled examples.
+In contrast, conditional entropy minimization [34] encourages the model to make confident predictions on unlabeled samples by minimizing the entropy of their predicted class distributions.
+Although this does not ensure correctness due to the absence of true labels, it promotes decisiveness and reduces model uncertainty.
+In this work, we adopt the strategy of using cosine similarity between the feature representations of labeled and unlabeled data as a criterion for pseudo-labeling of the unlabeled samples.
+We refer to [6] for a comprehensive review on the topic.
 2.3 Multimodality
 Our experience of the world is inherently multimodal we perceive objects through vision, hear sounds, feel textures, smell odors, and taste flavors.
-To develop an AI agent with similar capabilities, researchers have explored various methods Donato et al., Babadian et al..
-The simplest method to compensate multiple domains is using the concept of joint representation where the unimodal signals are combined in the same representation space by simple concatenation Baltrušaitis et al..
-Ngiam et al.
-used autoencoders for multimodal domains.
-They employed stacked denoising autoencoders to represent each modality individually and subsequently fusing them into a unified multimodal representation using an additional autoencoder layer.
-Similarly, Silberer and Lapata suggested a multimodal autoencoder for semantic concept grounding.
-They incorporated a reconstruction loss for training the representation and added a term in the loss function to predict object labels using the representation.
-Srivastava and Salakhutdinov introduced multimodal deep belief networks (DBNs) and multimodal deep boltzmann machines (DBMs) Srivastava and Salakhutdinov as multimodal representations.
-Kim et al.
-used a deep belief network for each modality and then combined them into joint representation for audiovisual emotion recognition.
-Huang and Kingsbury used a similar model for audio-visual speech recognition (AVSR), and Wu and Shao used it for audio and skeleton joint based gesture recognition.
-Ouyang et al.
-explored the use of multimodal DBMs for the task of human pose estimation from multi-view data.
-They demonstrated that integrating the data at a later stage after unimodal data, underwent nonlinear transformations, which was beneficial for the model.
-An alternative to a joint multimodal representation is a coordinated representation where instead of projecting the modalities together into a joint space, separate latent representations are learned for each modality and are coordinated through a constraint Frome et al..
-Sarfraz et al.
-recently utilized a multimodal CL dataset to learn complementary information from two different modalities.
-They employed the VGGSound dataset, which includes 10-second video clips along with their corresponding audio.
-On the other hand, Xu et al.
-developed an egocentric multimodal dataset using smart glasses equipped with sensors, combining data from cameras, accelerometers, and gyroscopes.
-This dataset covers 32 types of daily activities performed by 10 participants.
-They utilized the temporal binding network (TBN) Kazakos et al.
-to fuse the information from different sensors into single feature vector for further training and feature extraction.
-One of the major drawbacks of these methods is their inability to incrementally learn new domains.
-For example, if we train our model on two domains (sensor signals and images) and use a joint representation to learn these two modalities, introducing a new data modality later would necessitate re-training the model from scratch.
-Conversely, with our proposed exFeCAM model, we can add the new modality as a different domain and incrementally train the model without starting the training process anew.
-Additionally, our algorithm eliminates the necessity of storing data in memory for rehearsal, effectively addressing privacy concerns and storage challenges commonly encountered in exemplar-based CL methods.
-With recent advancements in large language models (LLMs) and the introduction of GPT-4 Achiam et al.
-and its variants, it is now possible to utilize the data coming from vision, audio, and text signals into a single model.
-However, these models face significant challenges, including high computational and memory demands and a lack of continual learning capabilities.
-Figure 1: Multimodal data comprising of two different domains, domain 1 contains the sensor information for different objects and domain 2 contains the images of those objects in different orientations, lightning conditions and different level of occlusion by the operator.
+To develop an AI agent with similar capabilities, several approaches have been explored [35], [36].
+The simplest method to compensate for multiple domains is using the concept of joint representation where the unimodal signals are combined in the same representation space by simple concatenation [37].
+[38] employed stacked denoising autoencoders to represent each modality individually and subsequently fusing them into a unified multimodal representation using an additional autoencoder layer.
+Similarly, [39] suggested a multimodal autoencoder for semantic concept grounding.
+They incorporated a reconstruction loss for training the representation and added a term in the loss function to predict object labels using the representation.[40] used a deep belief network (DBN) for each modality and then combined them into joint representations for audio-visual emotion recognition.
+A similar architecture was adopted by [41] for audio-visual speech recognition (AVSR), [42] applied the same strategy for integrating audio and skeletal joint representations in gesture recognition.
+[43] recently utilized a multimodal CL dataset to learn complementary information from two different modalities and performed a weighted fusion of the feature maps obtained from the respective feature extractors.
+On the other hand, [44] developed an egocentric multimodal dataset using smart glasses equipped with sensors and utilized the temporal binding network (TBN) [45] to fuse information from multiple sensors into a unified feature vector for feature extraction and training.
+However, a key limitation of these methods is their inability to incrementally adapt to new domains.
+Specifically, when training on multiple domains (such as sensor signals, visual images), the addition of a new modality often requires re-training the entire model from scratch.
+In contrast, our proposed exFeCAM algorithm addresses this challenge by treating each modality as a distinct domain, allowing for the incremental integration of new modalities without the need to restart training.
+Furthermore, our approach eliminates the need to store data in memory for rehearsal, effectively addressing privacy concerns and the storage constraints commonly associated with exemplar-based CL methods.
+Figure 1: Multimodal data comprising of two different domains, domain 1 contains the images of various objects in different orientations, lightning conditions, and different levels of occlusion by the operator and domain 2 contains the tactile sensor information for the corresponding objects.
 Moving from left to right the model trains in a class-incremental manner, while going from top-bottom it trains in a domain-incremental manner.
 3 Experimental Setup
-In this section, we describe the steps involved in the creation of the multimodal continual learning dataset as well as the standard evaluation criteria for the class incremental as well as domain incremental learning scenarios.
+In this section, we describe the steps involved in the creation of the multimodal continual learning dataset as well as the standard evaluation criteria for the class-incremental as well as domain-incremental learning scenarios.
 3.1 Continual Learning Scenario
-According to the division of incremental batches and the availability of task identities, continual learning can be broadly classified into three primary types or scenarios Van de Ven et al.:
-  • Task-Incremental Learning (TIL): It requires an algorithm to learn a set of distinct tasks progressively. Task identities/numbers are provided in both training and testing phase.
-  • Class-Incremental Learning (CIL): The algorithm must learn to differentiate between a growing number of classes incrementally. Task identities are only provided during training phase.
-  • Domain-Incremental Learning (DIL): The algorithm needs to learn the same tasks but in different domains. Task identities are not required.
-In addition to these major scenarios, there are also other scenarios present in the literature like task-free continual learning (TFCL), online continual learning (OCL), blurred boundary continual learning (BBCL), continual pre-training (CPT), we refer to Wang et al.
-for a thorough study.
-Figure 2: Commercial two fingers pneumatic gripper equipped with two force sensors and two flex sensors affixed together using stretchable nylon thread.
+According to the division of incremental batches and the availability of task identities, continual learning can be broadly classified into three primary types or scenarios [46]:
+  • Task-Incremental Learning (TIL): It requires an algorithm to learn a set of distinct tasks progressively, and it always knows which task it is working on.
+  • Class-Incremental Learning (CIL): The algorithm must learn to differentiate between a growing number of classes incrementally.
+  • Domain-Incremental Learning (DIL): The algorithm needs to learn the same task, but in different a domain or environment.
+In addition to these major scenarios, there are also other scenarios present in the literature like task-free continual learning (TFCL), online continual learning (OCL), blurred boundary continual learning (BBCL), continual pre-training (CPT), we refer to [5] for a thorough study.
+In our experiment, we designed a hybrid scenario that integrates both class-incremental and domain-incremental learning paradigms.
+Figure 2: Commercial two-finger pneumatic gripper equipped with two force sensors and two flex sensors affixed together using stretchable nylon thread.
 The flex sensor measures the bending of the finger and the force sensor measures the magnitude of the force applied to the fingertip.
 3.2 Multimodal Dataset
-The multimodal dataset comprises two components: sensor signals from an actuated finger and non-stationary images of objects captured by a vision camera.
-The two different modalities are fused into a hybrid setup that incorporates both class incremental as well as domain incremental scenarios as illustrated in Figure 1.
+The multimodal dataset comprises two components: sensor signal from an actuated finger and non-stationary images of objects captured by a vision camera.
+The two different modalities are structured into a CL setup that incorporates both class-incremental as well as domain-incremental scenarios as illustrated in Figure 1.
 In this setup, progression from left to right represents class-incremental training, while progression from top to bottom represents domain-incremental training.
 Both domains encompass five tasks, with each task involving the learning of two new classes.
-The first domain focuses on sensor signal data for various objects encountered in each experience and follows a class-incremental learning approach.
-The second domain deals with visual data of the same objects and employs both domain-incremental and class-incremental learning methods.
+The first domain focuses on visual data for various objects encountered in each experience and follows a class-incremental learning approach.
+The second domain deals with sensor signal data of the same objects and employs both domain-incremental and class-incremental learning methods.
 Further details on data collection for the different modalities are provided in the subsequent sections.
+In addition to our custom dataset, we also evaluate the effectiveness of our proposed method on a subset of the publicly available VGGSound dataset.
+We used this dataset due to it’s multimodal capabilities and it’s suitability for SSL settings.
+Specifically, we select five high-level parent classes: home, music, people, sports, and vehicles, each containing two specific subclasses.
+Every subclass includes 15 unique youtube videos for training, with each video lasting 240 seconds.
+Additionally, five separate videos per subclass are sampled in a similar fashion for testing.
+From these videos, we extract both the visual frames and the corresponding audio signals.
+To ensure accurate alignment between visual and audio modalities, we adopt the synchronization procedure outlined in [43].
+Finally, the extracted visual frames are resized to 32x32 pixels to reduce computational complexity and accelerate the training process.
 3.2.1 Sensor Signal
-To collect the sensor signals for the different objects we utilize a soft pneumatic gripper Hao et al.
-equipped with four sensors which are affixed together using stretchable nylon thread as illustrated in Figure 10.
-We have employed the commercially available flex sensors Saggio et al., Mishra et al.
-to measure the bending of the finger and a force sensitive resistor (FSR) Hollinger and Wanderley to quantify the force applied to the fingertip as shown in Figure 2.
-The outputs from the flex and force sensors are directed to an Arduino Due board, which in turn is connected to a computer for further processing and analysis.
-We gather the sensor data for 10 distinct objects, for each object the gripper acquires 50 data points, considering various orientations of the object.
-During the data collection process, for each data point the gripper holds and releases the object repeatedly at different contact points, with each cycle lasting $\sim$ 39 seconds.
-Each data point is represented as a 600-dimensional vector, encapsulating the sensor signals recorded during that time interval.
-The dataset acquisition for every object takes about 37 minutes to complete on a computer equipped with an Intel Core i7 CPU @2.80GHz (GTX 1070 GPU) running Windows 10.
+To collect sensor signals corresponding to different objects, we employ a soft pneumatic gripper [47] equipped with four sensors, which are secured in place using stretchable nylon thread.
+For capturing the bending motion of the fingers, we utilize commercially available flex sensors [48], [49], while fingertip force measurements are obtained using force-sensitive resistors (FSRs) [50], as illustrated in Figure 2.
+The outputs generated by both the flex and force sensors are routed to an Arduino Due board, which interfaces with a computer for subsequent signal processing and data analysis.
+We collect sensor readings across 10 different object classes, with each object contributing 50 individual data points that reflect variations in object orientation.
+During the data acquisition phase, the gripper holds and releases each object multiple times at different contact positions, with each cycle spanning approximately 39 seconds.
+Every data point is represented as a 600-dimensional vector that encapsulates the full set of sensor signals recorded within that duration.
+The complete data acquisition process for each object requires approximately 37 minutes and is carried out on a workstation equipped with an Intel Core i7 CPU @2.80GHz and a (GTX 1070 GPU).
 3.2.2 Image Data
@@
 The images are further resized to 32x32 dimension to speed up the training and evaluation process.
-3.3 Evaluation Metrics
-The performance of the CL algorithm depends on its ability to adapt, retain, and generalize the knowledge over time.
+3.3 Evaluation Metrics and Benchmark Algorithms
+The performance of the CL algorithm depends on it’s ability to adapt, retain, and generalize the knowledge over time.
 We evaluate the incremental accuracy (denoted as $a_{t}$) of the model after every task as well as the average accuracy over all the tasks for each domain denoted as $A_{T}$.
@@
 We define the expressions as follows:
-$A_{T}=\frac{1}{T}\sum_{t=1}^{t=T}a_{t}\\ A_{D}=\frac{1}{D}\sum_{T=1}^{T=D}A_{T}$ (1)
+$A_{T}=\frac{1}{T}\sum_{t=1}^{t=T}a_{t}\hskip 28.45274ptA_{D}=\frac{1}{D}\sum_{T=1}^{T=D}A_{T}$ (1)
 where $T$ and $D$ refer to the total number of incremental tasks in each domain and the total number of domains.
-To evaluate the SSL capabilities of the algorithm we perform two different experiments as proposed in the paper Smith et al., we use the following terminology to describe them: First, Random Images, where we have 20% of unlabeled data and 80% labeled data for each class.
-To collect the unsupervised data we randomly sample 20% (about 240) data from the training dataset for each class, remove the class label, shuffle, and store it in a separate folder.
+Additionally, to evaluate the SSL capability of our algorithm we perform two separate experiments as proposed in the paper [51], we use the following terminology to describe them: First, Random Images, where we have 30% of unlabeled data and 70% labeled data for each class.
+The unlabeled data is obtained by randomly sampling instances from the training dataset (ensuring that these images are removed from the labeled training set), removing their class labels, shuffling them, and storing them separately.
 Second, Unseen Images, where we sample the images of an unseen object of the same parent class from the training dataset and repeat the above steps.
-A point to note is that in the first case, the model might have already seen all the objects for each class, however for the second case, the model has never seen the objects present in the unsupervised dataset and thus makes it a more challenging situation.
-In addition, we also perform an ablation study to show the effectiveness of the intra-layer feature representation as well as SSL on the model’s overall accuracy.
-Table 1: Pre-train model architecture for both the domains of the custom multimodal dataset, domain 2 represents the image data and domain 1 signifies the sensor data.
+A point to note is that in the first case, the model might have already seen all the objects for each class however, for the second case, the model has never seen the objects present in the unsupervised dataset which makes it a more challenging situation.
+To assess the effectiveness of the proposed exFeCAM algorithm, we compare it against five benchmark methods: naive-concat, naive-add, vanilla FeCAM, FeTrIL, and a joint training approach.
+The naive-concat method employs a non-CL strategy by concatenating feature maps from both modalities for each class, followed by incremental training on newly introduced objects.
+In contrast, naive-add applies the same incremental learning procedure but replaces concatenation with a weighted summation of the feature maps.
+Both naive-concat and naive-add utilize an identical 3-layer feedforward neural network architecture, but without any CL capability.
+Vanilla FeCAM corresponds to the original FeCAM architecture introduced by [8].
+FeTrIL [52] also follows a replay free approach and stores the centroid representations of the past classes to produce their pseudo features by using geometric translation when training on new classes/domains.
+Lastly, the joint training approach involves offline training on the entire dataset, establishing an upper bound on accuracy.
+Figure 3: exFeCAM algorithm with intra-layer feature representation, online batchwise learning, and SSL capabilities for multimodal training.
+During training on multiple domains, the algorithm stores the mean and covariance matrices of the classes in a dictionary.
+In the testing phase, the algorithm uses these stored prototypes and employs the Mahalanobis distance to determine the prediction class.
+Supervised training is depicted with solid black lines, while unsupervised training is illustrated with dashed black lines.
+4 Methodology
+In this section, we discuss the modifications made to the FeCAM algorithm [8].
+Originally, the algorithm models the feature distribution of each class using a multivariate normal distribution, $\mathcal{N}(\mu,\Sigma)$, and employs a Mahalanobis distance based metric for classifying the respective samples.
+Although the authors proposed several techniques (co-variance matrix approximation, shrinkage and normalization) to enhance and stabilize the training process, one significant limitation of FeCAM is it’s requirement to store batches in temporary memory and conduct training only once all data for a particular task becomes available.
+This makes it unsuitable for online learning.
+To address this issue and improve the algorithm’s capability for multimodal training, as well as to adapt to the semi-supervised nature of the dataset, we introduce several new features to the exFeCAM algorithm, which are detailed below.
+4.1 Online Batchwise Training
+Online training refers to incrementally updating a model in a batch-wise fashion, rather than processing the entire dataset for a given task at once.
+In this approach, each batch of data is passed through a feature extractor to obtain feature maps, which are then used to calculate the prototypes (mean and co-variance matrix) for each class in that batch.
+A dictionary is initialized with class labels as keys and their corresponding prototypes as values.
+There are two possible scenarios when updating this dictionary:
+  1. 1.
+    If a class is encountered for the first time, a new entry is added with the class label as the key and the corresponding prototypes as the value.
+  2. 2.
+    If the class has been seen before, the existing prototypes for that class are retrieved and updated by averaging the previous prototypes with the new ones from the current batch.
+This method avoids the need to store incoming data for each task, which is critical for CL objectives and is especially suitable for autonomous agents with limited cache memory.
+Learning prototypes in a batch-wise manner addresses the challenge of not having access to previous data points, making it a more practical solution for real-time applications.
+Table 1: Pre-train model architecture for both the domains of the custom multimodal dataset, domain 1 represents the image data and domain 2 signifies the sensor data.
 The linear layers for both the models are removed once the pre-training is completed .
-Here, "classes" refers to the total number of output classes during the pre-training phase.
+Here, ”classes” refers to the total number of output classes during the pre-training phase.
 Model type Layer type Activation function Output channel/neurons Batch Normalization Other details
@@
 Conv2D Relu 64 Batchnorm2D kernel:3, stride:3, padding:0
-Domain 2 Conv2D Relu 128 Batchnorm2D kernel:2, stride:2, padding:0
+Domain 1 Conv2D Relu 128 Batchnorm2D kernel:2, stride:2, padding:0
 Flatten - - - -
@@
 Conv1D Leakyrelu (0.01) 128 Batchnorm1D kernel:3, stride:2, padding:1
-Domain 1 Flatten - - - -
+Domain 2 Flatten - - - -
 Linear Relu 256 - -
 Linear - classes - -
-4 Methodology
-In this section, we discuss the modifications made to the FeCAM algorithm Goswami et al..
-Originally, the algorithm models the feature distribution of classes using a multivariate normal distribution, $\mathcal{N}(\mu,\Sigma)$, and uses the Mahalanobis distance to perform classification.
-While the authors proposed several techniques to enhance and stabilize Mahalanobis-based distance classification, one significant limitation of FeCAM is its requirement to store batches in temporary memory and conduct training only once all data for a particular task is available.
-This makes it unsuitable for online learning.
-To address this issue and improve the algorithm’s capability for multimodal training, as well as to adapt to the semi-supervised nature of the dataset, we introduce several new features to the exFeCAM algorithm, which are detailed below.
-4.1 Online Batchwise Training
-Online training involves training a model incrementally in a batch-wise manner rather than on the entire dataset for a particular task at once.
-In this approach, each batch of data is passed through a feature extractor to obtain feature maps, which are then used to calculate the prototypes (mean and co-variance matrix) for each class in that batch.
-A dictionary is initialized with class labels as keys and their corresponding prototypes as values.
-There are two possible scenarios when updating this dictionary:
-  1. 1.
-    If a class is encountered for the first time, a new entry is added with the class label as the key and the prototypes as the value.
-  2. 2.
-    If the class has been seen before, the existing prototypes for that class are retrieved and updated by averaging the previous prototypes with the new ones from the current batch.
-This method avoids the need to store incoming data for each task, which is critical for continual learning (CL) objectives and is especially suitable for autonomous agents with limited cache memory.
-Learning prototypes in a batch-wise manner addresses the challenge of not having access to previous data points, making it a more practical solution for real-time applications.
 Algorithm 1 exFeCAM Algorithm
-Input: Data stream $(D_{0}^{0},D_{0}^{1},...,D_{1}^{0},D_{1}^{1},...,D_{n}^{0:t})$, where "$t$" refers to the number of tasks, "$n$" refers to the number of domains.
+Input: Data stream $\left(D_{0}^{0},D_{0}^{1},\dots,D_{1}^{0},D_{1}^{1},\dots,D^{0:t}_{n}\right)$
 Models: Feature extractor $f_{e}(x;\theta_{e})$, exFeCAM model $\phi(x;\mu,\Sigma)$
-Variables: $k$ : number of layers of feature extractor, $t$: total tasks, $n$: number of domains, $T$: total classes; $b$: batch size; $l,j$ : iterative variables
-$(x_{i},y_{i})\in D_{train}^{Supervised}$, $(x_{i}^{{}^{\prime}})\in D_{train}^{Unsupervised}$ , $(x_{i}^{{}^{\prime\prime}},y_{i}^{{}^{\prime\prime}})\in D_{test}$
+Variables: $k$: number of layers in feature extractor, $t$: total tasks, $n$: number of domains, $T$: total classes, $b$: batch size, $l,j$: iterative variables
+   $(x_{i},y_{i})\in D_{\text{train}}^{\text{Supervised}}$, $(x_{i}^{{}^{\prime}})\in D_{\text{train}}^{\text{Unsupervised}}$, $(x_{i}^{{}^{\prime\prime}},y_{i}^{{}^{\prime\prime}})\in D_{\text{test}}$
+   $\mu,\Sigma\leftarrow\{\}$
+Training Phase:
+1:for each experience $E_{i}\in\{E_{0},E_{1},\dots,E_{t}\}$ do
+2:  Temporary Buffer $(\epsilon)$ $\leftarrow\{\}$ {Temporary memory buffer for feature matching}
+3:  for each class $y_{i}\in\{y_{0},y_{1},\dots,y_{T}\}$ do
+4:   for $j=0$ to $b$ do
+5:    $\psi_{ij}\leftarrow\large[{f_{e}}{(x_{i})}_{k-(l-1)},{f_{e}}{(x_{i})}_{k-(l-2)},\dots,f_{e}(x_{i})_{k}$ ] {Concatenate intra-layer features}
+6:    $\mu[i]\leftarrow\frac{1}{b}\sum\psi_{ij}$ {Mean matrix}
+7:    $\Sigma[i]\leftarrow\text{Covariance}(\psi_{ij})$ {Covariance matrix}
+8:    $\Sigma[i]_{S}\leftarrow\text{Shrinkage}(\Sigma[i])$
+9:    $\Sigma[i]_{N}\leftarrow\text{Normalization}(\Sigma[i])_{S}$
+10:    if $\epsilon$ not empty then
 […]
-Training Phase
-for $E_{i}\in[E_{0},E_{1},...,E_{t}]$ do $\triangleright$ Different tasks/experiences
-     Temporary Buffer $(\epsilon)\leftarrow\{\}$triangleright Temporary memory buffer for feature matching
-     for $y_{i}\in[y_{0},y_{1},...,y_{T}]$ do $\triangleright$ Classes in each task
-         for $j\leftarrow$ $b$ do
-              $\psi_{ij}\leftarrow$ $[f_{e}(x_{i})_{k-(l-1)},f_{e}(x_{i})_{k-(l-2)},...,f_{e}(x_{i})_{k}]$triangleright Concatenate intra layer feature
-              $\mu[i]\leftarrow\frac{1}{b}\sum\psi_{ij}$triangleright Mean matrix
-              $\Sigma[i]\leftarrow$ Covariance($\psi_{ij}$) $\triangleright$ Covariance matrix
-              $\Sigma[i]_{S}\leftarrow$ Shrinkage($\Sigma[i]$)
-              $\Sigma[i]_{N}\leftarrow$ Normalization($\Sigma[i]_{S}$)
-              if $\epsilon$ is not empty then
-                  similarity $\leftarrow$ cosine similarity(features $\in\epsilon,x_{i}^{{}^{\prime}}$) $\triangleright$ Compute cosine similarity
-                  if similarity $\geq$ threshold  then
-                       Update $\mu[i]$, $\Sigma[i]_{N}$triangleright Update the mean and cov matrix for each class
-                  end if
-              end if
-         end for
-     end for
-end for
-Test Phase
-for $x_{i}^{{}^{\prime\prime}}\in D_{test}$ do
-     $\phi(f_{e}(x_{i}^{{}^{\prime\prime}});\mu,\Sigma)\leftarrow(f_{e}(x)-\mu_{i})^{T}$ $(\Sigma_{i})_{N}^{-1}$ $(f_{e}(x)-\mu_{i})$triangleright Calculate the Mahalanobis distance
-     $y_{pred}\leftarrow$ argmin $\phi(f_{e}(x_{i}^{{}^{\prime\prime}});\mu,\Sigma)$ $\forall$ $y_{i}=0,1,...T$triangleright Assign the class with minimum Mahalanobis distance
-end for
+12:     if similarity $\geq$ threshold then
+13:      Update $\mu[i]_{S},\Sigma[i]_{N}${Update the mean and cov matrix for each class}
+14:     end if
+15:    end if
+16:   end for
+17:  end for
+18:end for
+Test Phase:
+1:for each $x_{i}^{{}^{\prime\prime}}\in D_{\text{test}}$ do
+2:  $\phi(f_{e}(x_{i}^{{}^{\prime\prime}});\mu,\Sigma)\leftarrow(f_{e}(x)-\mu_{i})^{T}(\sum_{i})_{N}^{-1}(f_{e}(x)-\mu_{i})$ {Calculate the Mahalanobis distance}
+3:  $y_{\text{pred}}\leftarrow\arg\min\phi(f_{e}(x_{i}^{{}^{\prime\prime}});\mu,\Sigma),\forall y_{i}\in\{0,1,\dots,T\}${Assign the class with minimum Mahalanobis distance}
+4:end for
 4.2 Intra Layer Feature Representation
@@
 We argue that enriching the last layer representations with hierarchical intra-layer features increases robustness to domain shifts and thus improves generalizability to downstream continual tasks.
-These extracted features can be the class embeddings of a transformer encoder Vaswani et al.
-or flattened feature maps of a ResNet encoder He et al..
-There are two intuitive ways to promote intra-layer feature representations into the model: averaging the representations from the last "k" layers of the network or concatenating the representations from the last k layers.
+These extracted features can be the class embeddings of a transformer encoder [53] or the flattened feature maps of a ResNet encoder [54].
+There are two intuitive ways to promote intra-layer feature representations into the model: averaging the representations from the last ”k” layers of the network or concatenating the representations from the last ”k” layers.
 The first approach requires the output dimensions of all layers to be identical to facilitate summation.
 In our use case, we opt for concatenation, as the outputs from different layers have varying dimensions.
-Concatenation also preserves the rich, multi-scale features, unlike summation, which can result in information loss due to the merging of complementary features captured by different layers, potentially reducing the richness of the learned representations.
-For the multimodal experiment, we pre-train the domain 1 feature extractor on gripper data of objects not present in the training dataset, while the domain 2 feature extractor is trained on the Core50 dataset.
+Concatenation also preserves the rich, multi-scale features, unlike summation, which may result in information loss due to the merging of complementary features captured by different layers, potentially reducing the richness of the learned representations.
+For the multimodal experiment, we pre-train the domain 1 feature extractor on gripper data of objects not present in the training dataset, while the domain 2 feature extractor is pre-trained on the Core50 dataset [55].
+For the VGGSound experiment, we use the ResNet18 architecture as the feature extractor, pre-trained on the ImageNet1k dataset [56].
 More details about the pre-trained model architecture are provided in Table 1.
-For the Core50 experiment, we use the ResNet18 architecture as the feature extractor, pre-trained on the ImageNet1k dataset Russakovsky et al..
+Figure 4: On the left: intra layer feature representation for the custom dataset and the VGGSound dataset, we compare the improvement in the accuracy of the FeCAM algorithm on addition of knowledge from the last as well the penultimate layers of the pre-trained feature extractors.
+On the right: the plot shows sensor data accuracy (in orange) with and without the addition of random gaussian noise, the solid line in black shows the improvement in the prediction accuracy for the different experiences after addition of the image domain.
 4.3 Semi-Supervised Learning
-Semi-supervised learning (SSL) denotes the capability of an algorithm to learn from supervised data and then improve the performance of the model by training it on the unsupervised part.
-It is very useful in a situation where we have limited supervised data but we can collect huge amounts of unsupervised data from the environment.
-This is an active area of research given that large, labeled datasets are expensive, but most applications have access to plentiful, cheap unlabeled data.
-In this paper, we use the concept of pseudo-labeling to organize the unsupervised data and provide them with dummy labels for further training.
-To perform pseudo labeling we randomly store the feature maps of each class for every task in a temporary memory buffer and then perform cosine similarity between the reference feature maps and the feature maps generated from the complete unsupervised dataset as defined in equation 2,
-$similarity=\frac{x_{1}.x_{2}}{||x_{1}||_{2}||x_{2}||_{2}}$ (2)
-where $x_{1}$ and $x_{2}$ refers to the reference feature map present in the temporary memory and the feature maps from the unsupervised dataset, ||.$||.||_{2}$ refers to the l2 norm of the two tensors.
-If the value of the cosine similarity is greater than a certain threshold, we then train the exFeCAM algorithm on the respective feature maps and its associated pseudo labels.
-These pseudo-labels are derived from the class labels used for feature matching.
-The prototype values for that class are subsequently updated using the averaging rule described in the previous section.
-To determine the threshold value, we employ the Optuna framework Akiba et al., which uses bayesian optimization to explore the hyperparameter space efficiently.
-Figure 3: Modified FeCAM algorithm with intra layer feature representation, online batchwise learning, and semi-supervised learning capabilities for multimodal training.
-During training on multiple domains, the algorithm stores the mean and covariance matrices of the classes in a dictionary.
-In the testing phase, the algorithm uses these stored prototypes and employs the Mahalanobis distance to determine the prediction class.
-Supervised training is depicted with solid black lines, while unsupervised training is illustrated with dashed black lines.
+We incorporate an SSL approach to enable our algorithm to leverage unlabeled data, thus enhancing it’s overall generalization and improving performance on both previously encountered and newly introduced objects.
+To integrate SSL capabilities into the exFeCAM algorithm, we adopt a pseudo-labeling strategy, which allows us to organize the unlabeled data by assigning them dummy labels for further training.
+The pseudo-labeling process begins by randomly storing few feature maps of each class for every task in a temporary memory buffer.
+We then compute the cosine similarity between these reference feature maps and the feature maps generated from incoming unlabeled data, using the formula defined in Equation 2.
+[…]
+Here, $x_{1}$ represents a reference feature map from the temporary memory buffer, while $x_{2}$ corresponds to the feature map extracted from the new unlabeled data.
+The notation $||.||_{2}$ indicates the L2 norm of the respective vectors.
+If the cosine similarity between the two feature maps exceeds a predefined threshold, the exFeCAM algorithm proceeds to train on the new feature map and it’s assigned pseudo-label.
+These pseudo-labels are derived from the class labels associated with the most similar reference feature used during feature matching.
+The corresponding prototype values for that class, specifically the mean and covariance matrices are then updated according to the averaging rule discussed in the previous section.
+To select an appropriate threshold for pseudo-labeling, we utilize the Optuna framework [57], which employs bayesian optimization to efficiently explore the hyperparameter space and identify an optimal value.
+Once the pseudo-labeling process for a particular task is complete, the temporary memory buffer is discarded.
+Table 2: Comparison of the accuracy of the exFeCAM algorithm and the benchmark methods across two different multimodal datasets.
+The joint method represents a scenario where the algorithm has access to the entire dataset at once.
+Dataset Joint Naive-add Naive-concat FeTrIL Vanilla FeCAM ExFeCAM
+Custom multimodal 94.65±1.29 43.43±1.59 45.16±1.68 62.10±1.86 65.30±0.98 70.50±1.79
+VGGSound 92.48±2.31 29.78±1.76 30.94±1.81 38.21±0.89 39.96±1.16 45.96±1.98
 4.4 Multimodal Training
-We extend the class-incremental continual learning to a more realistic semi-supervised continual learning (SSCL) setting, where data distributions reflect existing object class correlations between, and among, the labeled and unlabeled data distributions.
-For our experiment we use 70% labeled data and 30% unlabeled data for the two protocols (random and unseen) discussed above.
-At task n, we denote batches of labeled data as $X_{n}={(x_{b},y_{b}):b_{i}\in(1,...,b)|y_{b}\in\tau_{n}}$ and batches of unlabeled data as $X_{n}^{{}^{\prime}}={(x_{b}^{{}^{\prime}}):b_{i}\in(1,...,b)}$, where $b$ refers to the batch size.
-The goal in task $n$ is to learn a model which predicts object class labels for any query input over all classes seen in the current and previous tasks ($\tau_{0}\cup\tau_{1}\cup...\cup\tau_{n}$).
-During each training step the model receives different classes incrementally over all the experiences for each domain as shown in Figure 1.
-The two domains contain data from different modalities specifically domain 1 contains the sensor data whereas domain 2 contains the vision data.
-The data for each task goes through a feature extractor to derive the feature maps from the different layers of the network.
-The feature maps for both domains are standardized to a fixed dimension of 384 and normalized to fall within the 0-1 range before being fed into the subsequent stages of the algorithm.
-These feature maps are then used to incrementally train the exFecam algorithm.
-During training for each task in the second domain, a small subset of the feature maps of the object is stored in a temporary memory buffer to facilitate pseudo-labeling for semi-supervised learning, as previously discussed.
-The feature maps are used to compute the mean and covariance matrix for each batch during training.
-The model performs covariance matrix approximation, shrinkage, and normalization to stabilize the training process Goswami et al..
-During the test phase, the extracted feature maps from the test dataset are used to calculate the Mahalanobis distance for all seen classes, and the class with the minimum distance value is selected as the output class.
-More details about the implementation of the algorithm is provided in the pseudo code.
-Additionally, we also perform a real-time test for object classification based on the multimodal data using the ROS framework as discussed in section 6 of the paper.
-Figure 4: Intra layer feature representation for the custom dataset and the Core50 dataset, we compare the improvement in the accuracy of the FeCAM algorithm on addition of knowledge from the last as well the penultimate layers of the pre-trained feature extractors.
-The improvement in case of Core50 dataset is considerably more than the custom dataset due to the presence of more diverse information in case of image data.
-We further analyzed the intra-layer feature representation for both the custom dataset and the Core50 dataset by comparing the accuracy improvement of the exFeCAM algorithm when incorporating knowledge from the last and penultimate layers of pre-trained feature extractors.
-The Core50 dataset showed a significantly greater improvement in accuracy compared to the custom dataset.
-The difference is attributed to the more diverse information contained in the image data of the Core50 dataset.
-Figure 5: On the left: radar plot shows the accuracy of the two domains after incremental training on different tasks.
-The FeCAM algorithm learns the first domain in a class-incremental approach, while it learns the second domain in a combined domain-incremental and class-incremental manner.
-On the right: the plot shows domain 1 accuracy (in orange) with random gaussian noise added to it, the solid line in grey shows the percentage improvement in the overall accuracy of the experiences after addition of the second domain.
-The black lines show the error bars for the two cases.
+We extend the traditional class-incremental CL framework to a more practical and challenging semi-supervised continual learning (SSCL) scenario.
+In this setting, the data distributions are designed to mirror real-world correlations among object classes by incorporating both labeled and unlabeled samples.
+For our experiments, we use a data split consisting of 70% labeled and 30% unlabeled data, applied under two protocols (random and unseen) as described in the previous section.
+At each task n, we represent the batches of labeled data as
+$X_{n}=\{(x_{b},y_{b}):b_{i}\in(1,...,b),\,y_{b}\in\tau_{n}\}$
+and the batches of unlabeled data as
+$X_{n}^{{}^{\prime}}=\{x_{b}^{{}^{\prime}}:b_{i}\in(1,...,b)\},$
+where $b$ denotes the batch size and $\tau_{n}$ indicates the set of classes associated with task $n$.
+The main goal during each task is to train a model that can accurately classify any query instance drawn from the union of all classes encountered in the current and preceding tasks, i.e., $\tau_{0}\cup\tau_{1}\cup\dots\cup\tau_{n}$.
+As training progresses, the model encounters new classes incrementally within each domain, as illustrated in Figure 1, where each domain corresponds to a different data modality.
+For each task, input data are passed through a feature extractor that produces feature maps from various layers of the network.
+These feature maps are standardized to a fixed dimension and normalized to 0–1 range before being fed into subsequent stages of the pipeline.
+These processed feature representations are then used to incrementally train the exFeCAM algorithm.
+To support the semi-supervised aspect of learning, a small subset of feature maps corresponding to the objects in each task is stored temporarily in a memory buffer.
+This buffer is utilized for pseudo-labeling of unlabeled data, as discussed in earlier sections.
+During the training phase, the algorithm computes the mean and covariance matrix for each batch of data and apply various techniques such as covariance matrix approximation, shrinkage, and normalization to ensure robust and stable learning, as described in the paper [8].
+In the testing phase, feature maps for the test instances are extracted using the appropriate feature extractor, and classification is performed by computing the Mahalanobis distance between each test sample and the class prototypes learned during training.
+The class corresponding to the minimum distance is selected as the predicted label.
+Further details about the algorithm implementation are provided in the corresponding pseudo-code.
+Figure 5: On the left, the radar plot illustrates the accuracy of various algorithms after incremental training on each experience for the custom multimodal dataset.
+The accuracy per experience represents the average performance across two domains: visual and sensor signals.
+On the right, the radar plot presents the per-experience accuracy for the VGGSound dataset, where the two domains consist of visual and audio data.
 5 Results
-To assess the performance of our algorithm and the effectiveness of the proposed modifications, we evaluate the exFeCAM algorithm on a custom multimodal dataset and the standard Core50 dataset used in the literature.
-The Core50 dataset was chosen for its non-iid nature and its compatibility for self-supervised learning.
-We do not explicitly compare our algorithm with other state-of-the-art (SOTA) benchmarks because the focus of this paper is not to establish a new level of accuracy for object classification, but rather to demonstrate the effectiveness of CL in a real world multimodal training setting.
-Nevertheless, our results are based on improvements to the original FeCAM algorithm, indirectly providing a comparison.
-All experiments are repeated three times, and we plot the mean and standard deviation of the respective metrics.
-Figure 6: Accuracy of the FeCAM algorithm on SSL tasks, the blue and green plot shows the SSL condition for unseen object from the same parent class, with blue indicating SSL enabled and green indicating SSL disabled case.
-The maroon and red plots represent the SSL condition for randomly selected images of the object, with maroon indicating SSL enabled and red indicating SSL disabled case.
-Top plot shows the accuracy for the custom multimodal dataset and the botton plot shows the accuracy for the Core50 dataset.
-The black dotted lines denotes the average accuarcy over all the experiences for different SSL cases.
-[…]
-[…]
-[…]
+To evaluate the performance of our algorithm and the effectiveness of the proposed modifications, we assess the exFeCAM algorithm on both a custom multimodal dataset and the VGGSound dataset.
+It is important to note that the primary focus of this paper is not to establish a new level of accuracy for object classification, but rather to demonstrate the effectiveness of CL in a challenging multimodal training setting.
+All experiments are conducted three times, and we report the mean and standard deviation of the relevant metrics.
+Table 2 compares the accuracy of various algorithms across the two multimodal datasets.
+Our proposed method achieves a significant improvement over all benchmark algorithms, which can be attributed to the integration of SSL and intra-layer feature representation (ILFR) capabilities in our architecture.
+Additionally, the vanilla FeCAM algorithm produces competitive results, likely due to it’s offline training mechanism.
+In contrast, our algorithm employs batch-wise online training, making it more relevant for real-world scenarios.
+FeTrIL algorithm also achieves a comparable accuracy, mostly attributed due to the presence of the pseudo-rehearsal technique to compensate forgetting of previous knowledge.
+Figure 4 (left) illustrates the enhancement in our algorithm’s prediction accuracy with the integration of intra-layer feature representation (ILFR).
+The accuracy is measured after the model has been incrementally trained on all tasks.
+For the custom dataset (represented in blue), the first five experiences depict the percentage improvement in domain 1 data (vision), while the subsequent five experiences reflect the accuracy gains in domain 2 (tactile sensor).
+Each experience consists of two distinct classes for both domains.
+On average, incorporating ILFR results in an overall accuracy improvement of approximately 3.12 $\pm$ 1.09 compared to the baseline without ILFR.
+A similar trend is observed in the VGGSound dataset (also shown in blue), where the initial five experiences correspond to image data, followed by five experiences with the associated audio data.
+In this case, the overall accuracy improvement is approximately 4.90 $\pm$ 1.20, further validating the benefits of ILFR.
+Figure 4 (right) demonstrates the advantage of training on a new domain when the information from the original domain (signals from the tactile sensor) is noisy.
+To simulate this scenario, gaussian noise is added to the sensor signal data.
+The solid orange line indicates the accuracy across different experiences for the signal domain with noisy sensor data, while the dashed orange line shows the original accuracy without noise.
+The solid black line indicates the improvement in accuracy when image data is incrementally introduced to mitigate the performance degradation caused by noisy sensor signals.
+The CL capabilities of exFeCAM, along with the benchmark algorithms, are presented in Figure 5.
+The spider plot depicts the combined accuracy for both domains, demonstrating that our proposed algorithm consistently outperforms all other methods across all experiences.
+The naive-concat and naive-add benchmark methods perform well in the initial experience but suffer significant accuracy degradation as the number of classes increases due to catastrophic forgetting.
+The joint method, which trains on all data at once, serves as an upper bound on per-experience accuracy.
+The FeTrIL algorithm initially shows results comparable to vanilla FeCAM, but it’s performance declines in later tasks, likely due to the moderate quality of the pseudo-rehearsal samples generated during training on new tasks.
+Figure 6: Accuracy of the exFeCAM algorithm across two different SSL scenarios.
+The blue plots correspond to the custom dataset, while the green plots represent the VGGSound dataset.
+Bar plots with solid lines indicate the condition where SSL is integrated into the algorithm, whereas dotted lines represent the condition without SSL.
+Additionally, we conducted a separate SSL experiment exclusively on the visual modality from both datasets, implementing two distinct types of experiments within the SSL protocol, as described in Section 3.3.
+Figure 6 presents the accuracy results for the custom dataset and the VGGSound dataset under both scenarios.
+The blue bar plot with a solid boundary represents the accuracy of the exFeCAM algorithm for the unseen object and random object scenarios with SSL enabled, while the dotted-boundary plot indicates the accuracy for the same scenarios without SSL.
+Similarly, the accuracy results for the VGGSound dataset are depicted using green bar plots.
+However, the overall accuracy of the VGGSound dataset is lower compared to the custom dataset, possibly due to the increased complexity and diversity of backgrounds in the extracted frames, whereas the custom dataset primarily consists of images with a relatively uniform background.
 Figure 7: The real-time multimodal experiment using ROS involves three publishers and one subscriber node.
@@
 The outputs from P1 and P2 are fed to the subscriber, where they undergo pre-processing, feature extraction using a pre-trained network, and finally to the FeCAM algorithm to make the predictions using the saved mean and co-variance matrices.
-Figure 4 (bottom) presents similar information for the Core50 dataset, where we employed the same experimental protocols.
-The key difference lies in the pre-trained model and the pre-training dataset used in this experiment.
-However, the improvement in accuracy for the Core50 dataset is minimal compared to the custom dataset.
-This could be attributed to the significant variations in the objects within each class, making the simple cosine similarity method less effective in this context.
-A more complex SSL method might be more beneficial for this scenario, which we plan to explore in future experiments.
 6 Real-Time Evaluation With ROS
-Robot Operating System (ROS) is a comprehensive framework that comprises a suite of tools, libraries, and protocols aimed at facilitating the development of various robotic systems.
+Robot Operating System (ROS) [9] is a comprehensive framework that comprises a suite of tools, libraries, and protocols aimed at facilitating the development of various robotic systems.
 The system manages the creation and control of communication between a robot’s peripheral modules, such as sensors, cameras, and actuators, thereby enabling the seamless integration of these components with parallization capabilties.
-Figure 8: Real-time images of the object from the camera’s perspective, along with the sensor data from four sensors attached to the soft finger, upon its actuation.
+Figure 8: Real-time images of the object from the camera’s perspective, along with the sensor data from four sensors attached to the soft finger, upon it’s actuation.
 These images and sensor readings are directly fed into the FeCAM model after resizing them to the desired dimension.
@@
 6.1 ROS Setup
-ROS is primarily designed to function optimally on Ubuntu, while only partially compatible with other operating systems, such as Windows and Mac.
-For our case we use ROS on Ubuntu 20.04 on a Intel Core i7 7th gen CPU @@@@ 2.80 GHz equipped with Nvidia GTX 1080 GPU.
-We use the Noetic Ninjemys distribution of ROS 1 Koubaa et al.
-in our experiment.
+ROS is primarily designed to function optimally on Ubuntu, while only partially compatible with other operating systems.
+For our case we use ROS on Ubuntu 20.04 on a Intel Core i7 7th gen CPU @@ 2.80 GHz equipped with Nvidia GTX 1080 GPU.
+We use the Noetic Ninjemys distribution of ROS 1 [9] in our experiment.
 A detailed explanation about the workflow of the project is provided in the flowchart in Figure 7.
-Figure 9: Square wave sent to the control box to actuate the soft pneumatic finger.
-We pressurise the finger for 3 seconds with a maximum pressure of 1.3 bars and then depressurise it for 2 seconds, the cycle is repeated "N" times till we reach the 15 seconds threshold for each run.
+Figure 9: Real-time object classification accuracy of the two modalities for four different objects.
+The grey plot represents the prediction accuracy in the image domain, while the pink plot highlights the accuracy for the sensor domain.
 The first publisher (P1) publishes the sensor signals coming from the four sensors attached to the soft finger, the second publisher (P2) publishes the video coming from the camera and the third publisher (P3) publishes the command pressure to the control box which inturn actuates the finger.
@@
 The data from P1 and P2 goes to the subscriber which then pre-process the data, extracts the frames from the video in case of P2 and sends it to the respective feature extractors.
-The output of the feature extractor is a 384-dimensional feature map (in case the intra-layer feature representation is turned on) which goes as an input to the exFecam model for further processing.
+The output of the feature extractor is a 384-dimensional feature map for both the domains which then goes as an input to the exFecam model for further processing and prediction.
 6.2 Real-Time Experiment
 To assess the performance of our algorithm, we conduct a quasi real-time experiment using a soft pneumatic gripper equipped with two flex sensors and two force sensors.
-During the evaluation phase, data from the sensors and the camera are collected in real-time for approximately 15 seconds.
-These data are treated as two distinct domains and are sent through the pipeline in parallel for further processing and subsequently to the exFecam model for predictions.
-To generate the predictions we utilize the stored mean and the co-variance matrix from a prior offline incremental experiment conducted for the two domains.
-Impressively, the mean and covariance matrix occupy only 22.5 MB of memory, making it quite efficient compared to other continual learning strategies
+During the evaluation phase, data from the tactile sensors and the camera are collected in real-time for approximately 15 seconds.
+These multimodal data are treated as two distinct domains and are sent through the dedicated pipeline in parallel for further processing and subsequently to the exFecam model for predictions as shown in Figure 7.
+To generate the predictions we utilize the stored mean and the co-variance matrix computed during the training phase.
+Impressively, the mean and covariance matrix occupy only 22.5 MB of memory, making it quite efficient compared to other CL strategies.
 However, since the model was initially trained on images of the objects in a different frame of reference, where the operator extends his arm and smoothly rotates the object in front of the camera.
@@
 To address this issue, for some objects, we occasionally need to refine the model’s knowledge by conducting a brief training session (lasting around 82 seconds) specifically for that object.
-Figure 8 shows a snapshot of the real-time data collection from both the sensors and the camera’s perspective.
+Figure 8 shows a snapshot of the real-time data collection from both the tactile sensors and the camera’s perspective.
 The image shows the progression of the sensor signal and the camera image upon actuating the finger.
 For each cycle the sensors collect 150 raw values, which are transformed into a single 600-dimensional vector, whereas the camera collects around 300 images of the object from various orientations within the 15-seconds interval.
-As discussed above one of the major advantages of this domain incremental training is for autonomous agents in situations where the field of view of one of the sensors gets restricted due to the presence of some obstacle from the environment, in that case the presence of other sensor can compensate for the missing knowledge.
+As discussed earlier, a significant advantage of domain-incremental CL training is it’s ability to incorporate new data modalities incrementally.
+This is particularly beneficial in scenarios where the input from one sensor becomes unreliable, either due to obstructions in the environment that limit it’s field of view or due to noise introduced by external factors.
+This advantage is further illustrated in Figure 9, where we evaluate object classification accuracy in real-time for four different objects: stapler, earpod, bottle, and marker pen.
+The results reveal that for the ”earpod” and ”marker pen”, the tactile sensor modality achieves notably higher classification accuracy than the corresponding visual modality.
+This discrepancy is likely caused by occlusions from the operator’s hand during image capture or limitations in the visual feature extractor’s ability to accurately represent small objects.
+In such cases, the availability of a second modality (here tactile sensor) effectively compensates for the missing or unreliable information, thereby improving overall classification performance.
+All the experiments were repeated 5 times for each of the four objects in real-time, and the final prediction accuracy was computed accordingly.
 6.3 Real World Application
-A real world application for such a technique can be in companion robots Broadbent et al., where the robot learns new tasks incrementally by interacting with its environment and recognizing different objects using information from multiple sensors available to it.
+A real world application for such a technique can be in companion robots [58], where the robot learns new tasks incrementally by interacting with it’s environment and recognizing different objects using information from multiple sensors available to it.
 The use of multiple sensors is particularly advantageous in scenarios where a single sensor cannot provide complete information about an object.
@@
 In such cases, information from a tactile sensor can be extremely beneficial.
-Our proposed framework simplifies the addition of new sensors by treating each as a new domain, eliminating the need to retrain the entire algorithm from scratch each time.
-All the sensors and the algorithm can be synchronized using a mobile version of ROS for embedded systems Min et al., providing a comprehensive solution.
+Our proposed framework simplifies the addition of new modalities by treating each as a new domain, eliminating the need to re-train the entire algorithm from scratch each time.
 7 Conclusion and Discussion
-In this paper, we present a novel continual learning method that combines the informations from different sensors in a domain incremental fashion while also learning new classes or objects in a class incremental manner.
-We enhanced the original FeCAM algorithm by incorporating online batch-wise training, intra-layer feature representation, and SSL capabilities thereby making the algorithm suitable for our scenario.
-The exFeCAM algorithm employs cosine similarity based feature matching to perform pseudo labeling when encountering unlabeled data.
-We validated the algorithm on two different modalities of data namely sensor signal based data and image based data, and demonstrated the ability of the algorithm to learn the two modalities incrementally.
-We also conducted an ablation study to show the contribution of different parts of the algorithm.
-Additionally, we performed a real-time experiment on a soft pneumatic gripper equipped with four sensors and an external camera system, all synchronised using the ROS framework.
-As part of our future work, we aim to enhance the exFeCAM algorithm further by adding forward and backward knowledge capabilities, either through interactions between class prototypes or by training a simple multilayer perceptron (MLP) network to learn these interactions.
-Moreover, currently we need to manually select the prediction domain modality in the presence of noise; in the future, we aim to enable the algorithm to automatically determine this by calculating the confidence of each prediction from the network.
-CRediT authorship contribution statement
-Nilay Kushawaha: Writing - original draft, Methodology, Investigation, Experimentation, Formal Analysis, Conceptualization. Egidio Falotico: Supervision, Funding acquisition, Project administration.
-Declaration of competing interest
+In this paper, we present a method for incrementally learning new data modalities without the need to re-train the model from scratch each time a new modality becomes available.
+Our approach combines class-incremental CL to progressively learn new object classes and domain-incremental CL to incorporate additional modalities for existing classes.
+To support this dual-incremental setup, we build upon the original FeCAM algorithm by introducing key enhancements, including online batch-wise training, intra-layer feature representation, and SSL capabilities, resulting in a more robust and versatile framework.
+Our method employs cosine similarity based feature matching to generate pseudo-labels for unlabeled data, enabling the model to effectively leverage both labeled and unlabeled examples during training.
+We evaluate the proposed algorithm on two multimodal datasets of varying complexity and demonstrate it’s ability to incrementally learn new classes in different domains.
+To better understand the contribution of individual components, we also conduct a comprehensive ablation study.
+To validate real-world applicability, we perform a real-time experiment using a soft pneumatic gripper equipped with four tactile sensors and an external camera.
+This experiment highlights the practical advantages of our multimodal CL approach in a realistic setting.
+As part of our future work, we plan to further improve the exFeCAM algorithm by incorporating forward and backward knowledge transfer, either through interactions between class prototypes or by training a multilayer perceptron (MLP) network to model these interactions.
+Additionally, instead of relying on a simple average of predictions from both modalities, we aim to adopt a more refined approach by using a weighted averaging scheme, where a separate network will estimate the confidence of each prediction to determine the appropriate weights.
+Acknowledgements
+We acknowledge the contribution from the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union–NextGenerationEU (Project IR0000011, CUP B51E22000150006, ”EBRAINS-Italy”)
+Conflicts of Interest
 The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
-Data availability
-Data will be made available on request. The code is available on the github repo : https://github.com/nilay121/Multimodal-Fusion-using-CL
-Acknowledgment
-We acknowledge the contribution from the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union–NextGenerationEU (Project IR0000011, CUP B51E22000150006, "EBRAINS-Italy")
+Data Availability Statement
+Data will be made available from the corresponding author upon reasonable request.
+Ethical Approval Statement
+All experiments were conducted solely by the author, and no other human participants were involved in the study.
+The author therefore gives informed written consent and confirms that no ethics committee approval is required, as per university guidelines.